### v1

# Paul Wild    Lutz Schröder    Karla Messing    Barbara König    Jonas Forster

Abstract
The classical Kantorovich-Rubinstein duality guarantees coincidence between metrics on the space of probability distributions defined on the one hand via transport plans (couplings) and on the other hand via price functions. Both constructions have been lifted to the level of generality of set functors, with the coupling-based construction referred to as the Wasserstein lifting, and the price-function-based construction as the Kantorovich lifting, both based on a choice of quantitative modalities for the given functor. It is known that every Wasserstein lifting can be expressed as a Kantorovich lifting; however, the latter in general needs to use additional modalities. We give an example showing that this cannot be avoided in general. We refer to cases in which the same modalities can be used as satisfying the generalized Kantorovich-Rubinstein duality. We establish the generalized Kantorovich-Rubinstein duality in this sense for two important cases: The Lévy-Prokhorov distance on distributions, which finds wide-spread applications in machine learning due to its favourable stability properties, and the standard metric on convex sets of distributions that arises by combining the Hausdorff and Kantorovich-Wasserstein distances.
1 Introduction
Measuring behavioural distances between probabilistic systems requires notions of distance between probability distributions (e.g. [29]). One well-established metric on the set of distributions over a metric space is variously termed the Kantorovich, Wasserstein, or Hutchinson metric. It can be calculated either by minimizing over the expected value of transport plans between, or couplings of, the given distributions, or by maximizing over the difference of expectations taken over all non-expansive price functions. The coincidence of these two values is the classical Kantorovich-Rubinstein duality [30, Theorem 5.10]. Intuitively speaking, a transport plan or coupling is a way to transform one distribution into another by shifting around weight, and its cost, to be minimized, is determined by how much weight is shifted over which distances. On the other hand, a price function determines a price for some commodity at given points; non-expansiveness of the price function means that no profit can be made from full-cost transport. The difference between the expected values of a price function under the given distributions is the profit to be made by having the commodity transported, and hence the amount one can offer to a logistics provider when outsourcing the transport.
Both the coupling-based definition and the price-function-based definition have been generalized categorically to construct liftings of set functors to the category of (pseudo-)metric spaces [4] and quantitative lax extensions of set functors [32], where the latter are distinguished by applying to unrestricted quantitative relations instead of only to (pseudo-)metrics. Metric functor liftings and lax extensions in particular serve to give a general treatment of behavioural distances on quantitative systems such as probabilistic, weighted, or metric [2] transition systems in the framework of universal coalgebra [25]. In this framework, functors serve as parameters determining a type of systems as their coalgebras; for instance, coalgebras for the distribution functor are Markov chains. In the generalized setting, the coupling-based construction is typically referred as the Wasserstein lifting or extension, and the price-function-based one as the Kantorovich lifting or extension. Both constructions are parametrized over a choice of quantitative modalities; the classical case involves, on both sides, only one modality, the expectation modality.
The interest in having a Kantorovich presentation of a given functor lifting or lax extension lies inter alia in the fact that one obtains a quantitative Hennessy-Milner property for the quantitative modal logic generated by the respective modalities [20, 32]. This property states coincidence of the behavioural distance induced by the given lifting or extension with the logical distance induced by the respective quantitative modal logic; thus, high distance between two states can always be certified by means of a modal formula, a principle that has received increased recent attention (e.g. [24, 28]). It has been shown that every metric functor lifting that preserves isometries [13] and every quantitative lax extension [32] is Kantorovich, i.e. can be presented via the generalized Kantorovich construction, using however a rather large (in particular typically infinite) set of modalities called the Moss modalities.
An important open point that remains is thus the question of what we term generalized Kantorovich-Rubinstein duality: In which cases does a Wasserstein distance given by a choice of modalities for a functor coincide with the Kantorovich distance for the same modalities? Known positive examples include, as mentioned, the classical Kantorovich-Wasserstein distance on distributions, but also the Hausdorff distance on the powerset. We begin our analysis by giving the first example of a natural Wasserstein metric for which generalized Kantorovich-Rubinstein duality in this sense fails, namely the standard $p$-Wasserstein metric for $p>1$ (which minimizes $p$-th roots of the expectation of $p$-th powers of couplings). As our main contribution, we then provide two new positive examples, namely the Lévy-Prokhorov distance on probability distributions [23] and the standard distance on the convex powerset, whose elements are convex sets of probability distributions. The Lévy-Prokhorov distance has seen a recent rise in popularity due to its favourable robustness properties that make it suitable for tasks in machine learning (such as conformal prediction [3] and corruption resistance [5]); moreover, it has been shown in recent work to induce precisely the behavioural distance defined by $\varepsilon$-bisimilarity [10]. In this case, the relevant modality is precisely the generally modality used in work on fuzzy description logics as an alternative formal correspondent of the natural-language term ‘probably’ [27]. The convex powerset plays a central role in the distribution semantics of Markov decision processes (or probabilistic automata) [7]. Its standard metric is just the composition of the Hausdorff and Kantorovich-Wasserstein metrics, and as such is again given via the Wasserstein construction for a modality composed from the expectation modality and the standard fuzzy diamond modality [32]. We show that the generalized Kantorovich-Rubinstein duality holds w.r.t. this modality. Beyond the mentioned benefits regarding characteristic quantitative modal logics, we demonstrate in this case that the new Kantorovich description also allows for more efficient computation of distances.
Related Work
Our Kantorovich presentation of convex powerset complements earlier results on a Wasserstein characterization [32] and a quantitative-algebraic presentation [21]. There has been recent work on what are termed correspondences between the generalized Kantorovich and Wasserstein constructions where a single modality is assumed for the Wasserstein construction, while an associated set of modalities is considered for the Kantorovich construction [16]. Correspondences in this sense thus lie between dualities as considered here, where we insist on the same modalities being used on both sides, and general theorems on Kantorovich presentations of lax extensions [32] and classes of functors [13] that use very large sets of liftings. Results are obtained for functors constructed from the main known instances (distributions with the standard Kantorovich-Wasserstein metric, powerset) by applying coproduct and product. The Kantorovich construction is sometimes referred to as the codensity construction [18, 19, 16], and as such has been used for logical characterizations of behavioural distances as mentioned above [19] but also for game characterizations [18]. The problem of generalized Kantorovich-Rubinstein duality has been stated already in work introducing coalgebraic behavioural distances [4], and a simple counterexample has been given; the counterexample we give here is distinguished by involving a quantitative modality that satisfies an analogue of two-valued separation [22].
2 Preliminaries
We discuss preliminaries on (pseudo-)metric spaces and coalgebras. Generally, we assume basic familiarity with category theory [1].
Metric spaces
We write $\oplus,\ominus\colon[0,1]\times[0,1]\to[0,1]$ for truncated addition and subtraction, i.e. $a\oplus b=\min\{1,a+b\}$ and $a\ominus b=\max\{0,a-b\}$. A ($1$-bounded) pseudometric space is a pair $(X,d_{X})$, where $X$ is a set and $d_{X}\colon X\times X\to[0,1]$ is a function, which for all $x,y,z\in X$ is subject to the conditions of reflexivity $d_{X}(x,x)=0$, triangle inequality $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$ and symmetry, that is $d_{X}(x,y)=d_{X}(y,x)$. A metric space is then a pseudometric space which is separated: If $d_{X}(x,y)=0$ then $x=y$. The Euclidean distance $d_{e}(a,b)=|b-a|$ makes $[0,1]$ into a metric space.
A function between the underlying sets $f\colon X\to Y$ of two pseudometric spaces $(X,d_{X})$, $(Y,d_{Y})$ is nonexpansive if distances are not increased by $f$, explicitly if for all $x,y\in X$ it holds that $d_{Y}(f(x),f(y))\leq d_{X}(x,y)$. Pseudometric spaces and nonexpansive functions between them form a category $\mathbf{PMet}$. The full subcategories of $\mathbf{PMet}$ spanned by metric spaces is denoted by $\mathbf{Met}$.
Coalgebra
Our main results derive some of their interest from their relevance to behavioural distances in coalgebras. Generally, the framework of universal coalgebra [25] is based on abstracting state-based systems as $F$-coalgebras for s functor $F\colon\mathbf{C}\to\mathbf{C}$ on a category $\mathbf{C}$, with $F$ determining the type of the system. Specifically, an $F$-coalgebra is a pair $(C,\gamma)$ consisting of a $\mathbf{C}$ object $C$, thought of as an object of states, and a morphism $\gamma\colon C\to FC$ determining transitions from states to structured collections of successor states, with the structure determined by $F$. A homomorphism between $F$-coalgebras $(C,\gamma)$ and $(D,\delta)$ is a $\mathbf{C}$-morphism $h\colon C\to D$ such that $\delta\circ h=Fh\circ\gamma$.
We list some common functors that occur in the description of $F$ in the definition above.
Example 1.
  1. 1.
    The (covariant) powerset functor $\mathcal{P}\colon\mathbf{Set}\to\mathbf{Set}$ sends each set to its powerset. On functions, $\mathcal{P}$ acts by taking images: For $A\in\mathcal{P}X$ and $f\colon X\to Y$ we have $\mathcal{P}f(A)=f[A]$. Its coalgebras are precisely sets equipped with a binary relation, i.e. transition systems or Kripke frames.
  2. 2.
    The finitely supported probability distribution functor $\mathcal{D}\colon\mathbf{Set}\to\mathbf{Set}$ sends a set $X$ to the set
    $\mathcal{D}X=\{\mu\colon X\to[0,1]\mid\mu(x)=0\text{ for almost all }x\in X\text{ and }\sum_{x\in X}\mu(x)=1\}$
    On a function $f\colon X\to Y$ the functor $\mathcal{D}$ measures probabilities of preimages:
    $\mathcal{D}f(\mu)(y)=\textstyle\sum_{x\in f^{-1}(y)}\mu(x).$
    The coalgebras of $\mathcal{D}$ are precisely Markov chains.
3 Dual Characterizations of Metrics
A central question in the study of state-based systems at large is whether two states exhibit the same behaviour. In universal coalgebra, answers for this type of question are provided by such concepts as Aczel-Mendler bisimulation or behavioural equivalence. When the behaviour of states has quantitative aspects, however, such as probabilistic transitions or outputs in a metric space, small deviations in these quantities immediately render two states behaviourally distinct under such two-valued notions. When one prefers to retain the information that these states differ only slightly, an established approach, discussed next, is to switch from behavioural equivalence relations to the more robust concept of behavioural metrics, equipping the state space with a pseudometric structure to describe how dissimilar individual states are in their behaviour.
A central role in the general coalgebraic treatment of behavioural distances is played by the concept of a functor lifting.
Definition 1.
Let $F\colon\mathbf{Set}\to\mathbf{Set}$ and $U\colon\mathbf{C}\to\mathbf{Set}$ be functors. A lifting of $F$ along $U$ is a functor $\overline{F}\colon\mathbf{C}\to\mathbf{C}$ such that the following diagram commutes.
$\mathbf{C}$$\mathbf{C}$$\mathbf{Set}$$\mathbf{Set}$$\scriptstyle{U}$$\scriptstyle{\overline{F}}$$\scriptstyle{U}$$\scriptstyle{F}$
When $\mathbf{C}$ is the category of pseudometric spaces and $U$ is the forgetful functor, i.e. the functor that maps pseudometric spaces to their underlying sets, the fibres above any set $X$ (the collection of pseudometric spaces carried by $X$) form a complete lattice under the pointwise order; we denote this lattice by $\mathbf{C}_{X}$. Given an $F$-coalgebra $(X,\gamma)$ and a functor lifting $\overline{F}$ we can construct a monotone function $\Phi_{\gamma}\colon\mathbf{C}_{X}\to\mathbf{C}_{X}$ on this complete lattice, sending a pseudometric $d_{X}\colon X\times X\to[0,1]$ to $\Phi_{\gamma}(d_{X})$ given by
$\Phi_{\gamma}(d_{X})(x,y)=d_{\overline{F}(X,d_{X})}(\gamma(x),\gamma(y)).$
The behavioural distance [4] on $(X,\gamma)$ is then defined as the least fixpoint $\mu\Phi_{\gamma}$ of the function $\Phi_{\gamma}$, which exists by the Knaster-Tarski fixpoint theorem.
Example 2.
  1. 1.
    The Hausdorff lifting $\overline{\mathcal{P}}\colon\mathbf{PMet}\to\mathbf{PMet}$ equips the powerset $\mathcal{P}X$ of the carrier of a metric space $(X,d_{X})$ with the Hausdorff metric $\delta^{\mathsf{H}}(d_{X})$. The distance of two subsets $A,B\in\mathcal{P}X$ is then given by
    $\delta^{\mathsf{H}}(d_{X})(A,B):=\max(\adjustlimits{\sup}_{x\in A}{\inf}_{y\in B}d_{X}(x,y),\adjustlimits{\sup}_{y\in B}{\inf}_{x\in A}d_{X}(x,y))$
  2. 2.
    The Kantorovich-Wasserstein lifting $\overline{\mathcal{D}}\colon\mathbf{PMet}\to\mathbf{PMet}$ of $\mathcal{D}$ equips the set $\mathcal{D}X$ with the Kantorovich-Wasserstein distance $\delta^{\mathsf{K}}(d_{X})$, which is defined as
    $\delta^{\mathsf{K}}(d_{X})(\mu,\nu):=\sup\{\mathbb{E}_{\nu}({f})-\mathbb{E}_{\mu}({f})\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexpansive}\},$
    where $\mathbb{E}_{\mu}({f})=\sum_{x\in X}\mu(x)\cdot f(x)$ denotes the expected value of $f$ under $\mu$.
The examples above can be seen as instances of more general constructions introduced below, which are parametric in a $\mathbf{Set}$-endofunctor and a set of $[0,1]$-valued predicate liftings.
Definition 2.
  1. 1.
    A $[0,1]$-valued predicate lifting for a functor $F\colon\mathbf{Set}\to\mathbf{Set}$ is a natural transformation of type $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$. A predicate lifting $\lambda$ is well-behaved if the following conditions hold:
    • Monotonicity: If $f\leq g$, then $\lambda_{X}(f)\leq\lambda_{X}(g)$, where the ordering on functions is computed pointwise.
    • Subadditivity: for $f,g\in[0,1]^{X}$, we have $\lambda_{X}(f\oplus g)\leq\lambda_{X}(f)\oplus\lambda_{X}(g)$, where the sum of two functions is calculated pointwise.
    • Zero preservation: $\lambda_{X}(0_{X})=0_{FX}$, where $0_{X}$, $0_{FX}$ are the constant zero functions on the respective sets.
  2. 2.
    Let $\lambda$ be a predicate lifting for $F$. The Kantorovich lifting of $F$ to the category of pseudometric spaces sends a metric $d_{X}$ to $K_{\lambda}(d_{X})\colon FX\times FX\to[0,1]$ defined by
    $K_{\lambda}(d_{X})(s,t):=\sup\{|\lambda_{X}(f)(t)-\lambda_{X}(f)(s)|\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexp.}\}$
    If $\Lambda$ is a set of predicate liftings, we put $K_{\Lambda}=\sup_{\lambda\in\Lambda}K_{\lambda}$ (pointwise).
  3. 3.
    Let $s,t\in FX$. The set of couplings $\Gamma(s,t)$ is defined as
    $\Gamma(s,t):=\{c\in F(X\times X)\mid F\pi_{1}(c)=s\text{ and }F\pi_{2}(c)=t\}$
  4. 4.
    Let $\lambda$ be a well-behaved predicate lifting for $F$ and assume that $F$ preserves weak pullbacks. The Wasserstein lifting of $F$ to the category of pseudometric spaces is defined as the lifting that equips $FX$ with $W_{\lambda}(d_{X})$ where
    $W_{\lambda}(d_{X})(s,t)=\inf\{\lambda_{X\times X}(d_{X})(c)\mid c\in\Gamma(s,t)\}.$
While the Kantorovich lifting assumes no conditions on supplied structures, the Wasserstein lifting is significantly more particular, requiring both pullback preservation of the underlying functor and that the predicate lifting be well-behaved, to ensure that $W_{\lambda}(d_{X})$ is a pseudometric whenever $d_{X}$ is [4, 6].
Remark 1.
It is well known that predicate liftings correspond to simple morphisms, sometimes dubbed evaluation functions, by the Yoneda lemma [26]. In the concrete instance of $[0,1]$-valued predicate liftings, we have that natural transformations of the form $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$ are in bijection with morphisms of type $\mathsf{ev}_{\lambda}\colon F[0,1]\to[0,1]$. Then the condition of a predicate lifting being well-behaved translates roughly to the corresponding evaluation function being well-behaved [4, 32].
Example 3.
  1. 1.
    Let $F=\mathcal{P}$ and $\lambda\colon[0,1]^{X}\to[0,1]^{\mathcal{P}-}$ be the natural transformation whose components calculate suprema of images: For $A\subseteq X$ and $f\in[0,1]^{X}$ we define $\lambda_{X}(f)(A)=\sup f[A]$. Then $\delta^{\mathsf{H}}=K_{\lambda}=W_{\lambda}$.
  2. 2.
    Let $F=\mathcal{D}$ and let $\lambda$ be the predicate lifting calculating expected values: For $\mu\in\mathcal{D}X$ and $f\in[0,1]^{X}$ we have $\lambda_{X}(f)(\mu)=\mathbb{E}_{\mu}({f})$. Then $\delta^{\mathsf{K}}=K_{\lambda}=W_{\lambda}$.
If, like in the two examples above, the categorical Kantorovich and Wasserstein constructions $K_{\lambda}$ and $W_{\lambda}$ coincide, we say that generalized Kantorovich-Rubinstein duality holds. This name is motivated by the particular case of Example˜3.2, the classical Kantorovich-Rubinstein duality that dates back to the beginnings of transportation theory [17]. It is important to note that generalized duality in this sense may fail, as demonstrated by the case of $p$-Wasserstein distance that we discuss later. It is a general fact [4, Theorem 5.27] that $K_{\lambda}\leq W_{\lambda}$, so duality hinges on the inequality $W_{\lambda}\leq K_{\lambda}$.
Transportation Theory
The Wasserstein presentation is closely related to optimal transportation theory: In the case of the Kantorovich-Wasserstein lifting, one may view probability distributions $\mu,\nu$ as producers and consumers of a resource respectively (with the amount of the resource produced/consumed being fixed to $1$). Now one wants to transport the produced resources to the consumers in the economically most efficient possible way, minimizing the average distance each unit of resource needs to travel. Couplings $c$ of $\mu$ and $\nu$ can then be seen as transport plans, with $\mathbb{E}_{c}({d_{X}})$ giving the total cost of carrying out the plan. Then the Kantorovich-Wasserstein distance gives us by definition the minimal possible cost.
In the distribution case (Example˜3.2), the Kantorovich presentation can be explained via the analogy of “outsourcing the transport” by defining a function $f$ that assigns a price to each $x\in X$. This function must satisfy the requirement that it is non-expansive, i.e., that the difference of prices assigned to $x,y$ is always at most $d_{X}(x,y)$ (meaning that no extra profit can be made from such a transport). The overall profit under such a price function is then the income obtained from the consumers ($\mathbb{E}_{\nu}({f})$) minus the cost paid to the producers ($\mathbb{E}_{\mu}({f})$). Taking the maximum over all such $f$ gives us the value of the Kantorovich lifting.
Expressive Logics
The Kantorovich lifting is closely related to characteristic multi-valued modal logics: These are logics in which formulae $\varphi$ receive semantics in coalgebras $(X,\gamma)$, inducing an interpretation function $\llbracket\varphi\rrbracket_{\gamma}\colon X\to[0,1]$. The logical distance of two states is then the supremum of all distances witnessed by such formulas.
The semantics of a modal operator $L$ in these types of logics is usually given as a predicate lifting $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$, with $\llbracket L\varphi\rrbracket_{\gamma}$ being inductively defined as $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$. Then expressivity of the logic (the fact that the behavioural distance can be witnessed by formulae of the logic arbitrarily closely) can be shown by exploiting the Kantorovich-presentation of the lifting. In fact, the interpretation $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$ can be viewed as emulating one step of the functional $\Phi_{\gamma}$, under the condition that the interpretations of formulae $\llbracket\varphi\rrbracket_{\gamma}$ are able to approximate any nonexpansive function $f\colon X\to[0,1]$ arbitrarily closely [12].
Hence, the duality can be used very fruitfully: One can use the Wasserstein view to compute (an under-approximation of) the behavioural distance and switch to Kantorovich to determine the distinguishing formula witnessing this distance.
The $p$-Wasserstein Distance
There also exists a parametrized version of the Kantorovich-Wasserstein distance, called the $p$-Wasserstein distance for some real parameter $p\geq 1$}. In categorical terms, it is given by the predicate lifting $\lambda_{p}(f)(\mu)=(\mathbb{E}_{\mu}({f^{p}}))^{\frac{1}{p}}$ (where $f^{p}$ takes the $p$-th power pointwise), that is:
$W_{\lambda_{p}}(d)(\mu,\nu)=\inf\{(\mathbb{E}_{\rho}({d^{p}}))^{\frac{1}{p}}\mid\rho\text{ is a coupling of $\mu$ and $\nu$}\}.$
For $p=1$ this is just the usual Kantorovich-Wasserstein distance and duality holds. If $p>1$, however, then duality may fail, and the corresponding Kantorovich construction $K_{\lambda_{p}}$ may be strictly below $W_{\lambda_{p}}$:
Example 4.
Let $p=2$. Let $(X,d)$ be a two-element discrete metric space, that is $X=\{0,1\}$ and $d(0,1)=d(1,0)=1$, and let $\mu=\frac{2}{3}\cdot 0+\frac{1}{3}\cdot 1$ and $\nu=\frac{1}{3}\cdot 0+\frac{2}{3}\cdot 1$. Then we have $K_{\lambda_{p}}(d)(\mu,\nu)\leq\frac{1}{3}<\frac{1}{\sqrt{3}}=W_{\lambda_{p}}(d)(\mu,\nu)$.
Proof.
We begin by showing the last equality. We note that $d^{2}=d$, so that
$W_{\lambda_{2}}(d)(\mu,\nu)=(\delta^{\mathsf{K}}(d)(\mu,\nu))^{\frac{1}{2}}=(\tfrac{1}{3})^{\frac{1}{2}}=\tfrac{1}{\sqrt{3}},$
where in the first step we used that $(-)^{\frac{1}{2}}$ is monotone and continuous.
For the first inequality, let $f\colon X\to[0,1]$ be nonexpansive and put $a=f(0)$ and $b=f(1)$. We show that $\lambda_{2}(f)(\nu)-\lambda_{2}(f)(\mu)\leq\frac{1}{3}$; the proof that $\lambda_{2}(f)(\mu)-\lambda_{2}(f)(\nu)\leq\frac{1}{3}$ is analogous. We may assume wlog. that $a<b$, as otherwise the left hand side of our target inequality is nonpositive. Define $g\colon[0,1]\to\real$ via $g(t)=\sqrt{(1-t)a^{2}+tb^{2}}$. Then we have $g(0)=a$, $g(\frac{1}{3})=\lambda_{2}(f)(\mu)$, $g(\frac{2}{3})=\lambda_{2}(f)(\nu)$ and $g(1)=b$. As the function $g$ is concave, we also have
$g(0)+g(1)=(\tfrac{1}{3}\cdot g(0)+\tfrac{2}{3}\cdot g(1))+(\tfrac{2}{3}\cdot g(0)+\tfrac{1}{3}\cdot g(1))\\ \leq g(\tfrac{1}{3}\cdot 0+\tfrac{2}{3}\cdot 1)+g(\tfrac{2}{3}\cdot 0+\tfrac{1}{3}\cdot 1)=g(\tfrac{1}{3})+g(\tfrac{2}{3})$ (1)
Therefore we have:
$(g(\tfrac{2}{3})-g(\tfrac{1}{3}))\cdot(g(\tfrac{2}{3})+g(\tfrac{1}{3}))$ $=g(\tfrac{2}{3})^{2}-g(\tfrac{1}{3})^{2}$
$=(\tfrac{1}{3}a^{2}+\tfrac{2}{3}b^{2})-(\tfrac{2}{3}a^{2}+\tfrac{1}{3}b^{2})$
$=\tfrac{1}{3}(b^{2}-a^{2})$
$=\tfrac{1}{3}(b-a)(a+b)$
$\leq\tfrac{1}{3}(g(0)+g(1))$ $(\text{$f$ is nonexpansive})$
[…]
Dividing by $g(\tfrac{2}{3})+g(\tfrac{1}{3})$ on both sides results in the claimed inequality. ∎
4 Lévy-Prokhorov Distance
The Lévy-Prokhorov distance provides an alternative to the Kantorovich-Wasserstein distance when it comes to measuring the distance between probability distributions. If $(X,d)$ is pseudometric space, and $\mu,\nu\in\mathcal{D}X$ are (discrete) probability distributions, then we define
$\delta^{\mathsf{LP}}(d)(\mu,\nu)=\inf\{\varepsilon\mid\forall A\subseteq X.\;\mu(A)\leq\nu(A^{d}_{\varepsilon})+\varepsilon\},$
where $A^{d}_{\varepsilon}=\{y\in Y\mid\inf_{x\in A}d(x,y)\leq\varepsilon\}$. The definition of the Lévy-Prokhorov distance sometimes includes the mirrored condition $\forall B\subseteq Y.\;\nu(B)\leq\mu(B^{d}_{\varepsilon})+\varepsilon$, but this second clause is redundant and does not actually change the induced pseudometric.
The Lévy-Prokhorov distance has recently been investigated by Desharnais and Sokolova [11], who prove that it is a functor lifting, but not a monad lifting, and that it characterizes the notion of $\varepsilon$-bisimulation [10].
The Lévy-Prokhorov distance admits a Wasserstein representation, which is sometimes also known as the Ky Fan metric. It is typically phrased in terms of pairs of (not necessarily independent) random variables that are distributed according to the given distributions, but this is equivalent to considering all their couplings (cf. Section˜2). The predicate lifting $\lambda$ underlying this representation is given by
$\lambda_{X}(f)(\mu)=\inf\{\varepsilon\geq 0\mid\mu(\{x\in X\mid f(x)>\varepsilon\})\leq\varepsilon\},$ (2)
and using this predicate lifting we have $\delta^{\mathsf{LP}}(d)=W_{\lambda}(d)$ for every pseudometric $d$, explicitly:
$\delta^{\mathsf{LP}}(d)(\mu,\nu)=\inf\{\inf\{\varepsilon\geq 0\mid\rho(\{(x,y)\in X\times Y\mid d(x,y)>\varepsilon\})\leq\varepsilon\}\\ \mid\rho\in\Gamma(\mu,\nu)\}.$
The predicate lifting from (2) has independently been used under the name ‘generally’ in work on fuzzy description logics [27], and it admits a number of equivalent representations. Intuitively, all of these representations amount to the statement that the value $\lambda_{X}(f)(\mu)$ is high if the value of $f$ is high with high probability.
Lemma 1.
lem:lp-pred-lifting Let $X$ be a set, let $f\colon X\to[0,1]$, and let $\mu\in\mathcal{D}X$. Then we have:
  1. 1.
    $\lambda_{X}(f)(\mu)=\inf_{\varepsilon\geq 0}\max(\mu(\{x\in X\mid f(x)>\varepsilon\}),\varepsilon)$
  2. 2.
    $\lambda_{X}(f)(\mu)=\sup_{\varepsilon\geq 0}\min(\mu(\{x\in X\mid f(x)>\varepsilon\}),\varepsilon)$
  3. 3.
    $\lambda_{X}(f)(\mu)=\sup\{\varepsilon\geq 0\mid\mu(\{x\in X\mid f(x)>\varepsilon\})\geq\varepsilon\}$
All these identities, and also (2), remain true if $f(x)>\varepsilon$ is replaced by $f(x)\geq\varepsilon$.
It follows from LABEL:lem:lp-pred-lifting that $\lambda$ is self-dual (equal to its own dual):
Lemma 2.
lem:lp-pred-lifting-self-dual For every $X$, $f$ and $\mu$ we have $\lambda_{X}(f)(\mu)=1-\lambda_{X}(1-f)(\mu)$.
The Wasserstein representation above is justified by the fact that the predicate lifting is well-behaved:
Lemma 3.
lem:lp-pred-lifting-wb The predicate lifting $\lambda$ as per (2) is well-behaved.
4.1 Duality
Next, we show that the Lévy-Prokhorov distance admits a dual representation using the same predicate lifting $\lambda$, that is, we have $K_{\lambda}=W_{\lambda}$. We prove this duality in the more general setting where the two constructions apply to fuzzy relations that need not be pseudometrics. Recall that a fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ between sets $X$ and $Y$ is a function $r\colon X\times Y\to[0,1]$. The Wasserstein construction applies to fuzzy relations the same way it does to pseudometrics, while the Kantorovich construction is defined in terms of pairs of functions that satisfy a nonexpansiveness condition with respect to the given fuzzy relation. They are therefore both examples of (fuzzy) relational liftings or relators (e.g. [14] and references therein), as they lift fuzzy relations of type $X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to relations of type $FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$:
Definition 3.
Let $\lambda$ be a monotone predicate lifting for a set functor $F$, and let $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$.
  1. 1.
    The relational Wasserstein lifting $W_{\lambda}(r)\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
    $W_{\lambda}(r)(s,t)=\inf\{\lambda_{X\times Y}(r)(c)\mid c\in\Gamma(s,t)\}$
    for every $s\in FX$ and $t\in FY$.
  2. 2.
    An $r$-nonexpansive pair is a pair of functions $(f,g)$ where $f\colon X\to[0,1]$, $g\colon Y\to[0,1]$ and $g(y)-f(x)\leq r(x,y)$ for all $x\in X$ and $y\in Y$.
  3. 3.
    The relational Kantorovich lifting $K^{\mathsf{rel}}_{\lambda}\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
    $K^{\mathsf{rel}}_{\lambda}(r)(s,t)=\sup\{\lambda_{Y}(g)(t)\ominus\lambda_{X}(f)(s)\mid(f,g)\text{ $r$-nonexpansive}\}$
    for every $s\in FX$ and $t\in FY$. Additionally, put $K^{\mathsf{rel}}_{\Lambda}=\sup_{\lambda\in\Lambda}K^{\mathsf{rel}}_{\lambda}$ if $\Lambda$ is a set of predicate liftings.
Both of these constructions satisfy certain laws (that we will not restate here) making them lax extensions. Wild and Schröder [32] give results that relate $K^{\mathsf{rel}}$ to its pseudometric counterpart. The most relevant consequence of these results for our purposes is the following:
Lemma 4.
lem:self-dual-rel-equals-sym If $\lambda$ is a self-dual predicate lifting, then $K^{\mathsf{rel}}_{\lambda}(d)=K_{\lambda}(d)$ for every pseudometric $d$.
Out of the two representations of the Lévy-Prokhorov distance discussed earlier, the second, being a Wasserstein lifting, readily generalizes to fuzzy relations. Therefore, we define the relational Lévy-Prokhorov lifting $\delta^{\mathsf{LP}}$ to be the assignment that maps each fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to $\delta^{\mathsf{LP}}(r)=W_{\lambda}(r)\colon\mathcal{D}X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}\mathcal{D}Y$.
We discussed in Section˜3 that the inequality $\leq$’ follows from the general theory of Kantorovich and Wasserstein liftings. The same is true for the respective lax extensions [31, Lemma 5.22], so that it suffices to prove the converse inequality $\geq$’. In the proof of the classical Kantorovich-Rubinstein duality (e.g [30, Theorem 5.10]), this direction amounts to, given an optimal transport plan in the shape of a coupling between the distributions at hand, constructing two price functions that correspond to the optimal cost, in the sense that they form a nonexpansive pair that witnesses the supremum in the definition of the relational Kantorovich lifting $K_{\mathbb{E}}$. In our proof of Lévy-Prokhorov duality we use a similar approach, which means that we should first understand how to phrase computation of the Lévy-Prokhorov distance in terms of a transport problem.
Let $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ and let $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$. The Wasserstein representation $\delta^{\mathsf{LP}}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu)$ can be rewritten for this purpose. For $\varepsilon\geq 0$, define $r^{\varepsilon}(x,y)=0$ if $r(x,y)<\varepsilon$ and $r^{\varepsilon}(x,y)=1$ otherwise. Then $\rho(\{(x,y)\mid r(x,y)\geq\varepsilon\})=\mathbb{E}_{\rho}({r^{\varepsilon}})$ for every $\rho\in\mathcal{D}(X\times Y)$, and therefore we have
$W_{\lambda}(r)(\mu,\nu)=\inf\{\inf_{\varepsilon\geq 0}\max(\varepsilon,\mathbb{E}_{\rho}({r^{\varepsilon}}))\mid\rho\in\Gamma(\mu,\nu)\}=\inf_{\varepsilon\geq 0}\max(\varepsilon,W_{\mathbb{E}}(r^{\varepsilon})(\mu,\nu)),$
where we used the representation of $\lambda$ from LABEL:lem:lp-pred-lifting.1 while replacing the strict inequality by a non-strict one and then swapped the infimum over couplings all the way to the inside.
This means that Lévy-Prokhorov distance is determined by the solutions to the transport problems for the $r^{\varepsilon}$. As each such $r^{\varepsilon}$ is a crisp relation (i.e. only has 0 and $1$ entries), the optimal price functions can be made crisp as well:
Lemma 5.
lem:01-price-functions Let $r\colon X\times Y\to\{0,1\}$ be a crisp relation and let $\mu\in\mathcal{D}X$, $\nu\in\mathcal{D}Y$. Then there exist functions $f\colon X\to\{0,1\}$ and $g\colon Y\to\{0,1\}$ such that $(f,g)$ is an $r$-nonexpansive pair and $\mathbb{E}_{\nu}({g})-\mathbb{E}_{\mu}({f})\geq W_{\mathbb{E}}(r)(\mu,\nu)$.
This allows us to establish duality:
Theorem 6.
thm:levy-prokhorov-duality-relational For every $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ and every $\mu\in\mathcal{D}X$, $\nu\in\mathcal{D}Y$,
$\delta^{\mathsf{LP}}(r)(\mu,\nu)=K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu).$
Proof(sketch).
As mentioned before, we only need to prove $W_{\lambda}(r)\leq K^{\mathsf{rel}}_{\lambda}(r)$. Assume $\varepsilon<W_{\lambda}(r)(\mu,\nu)$. Then there is an $r^{\varepsilon}$-nonexpansive pair $(p,q)$ of crisp price functions witnessing the transport cost wrt. $r^{\varepsilon}$. One then replaces the function values 0 and $1$ by $\mathbb{E}_{\mu}({p})$ and $\mathbb{E}_{\mu}({p})+\varepsilon$ to arrive at a $d$-nonexpansive pair $(f,g)$. Using this pair we show
$K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)\geq\lambda_{Y}(g)(\nu)-\lambda_{X}(f)(\mu)\geq\mathbb{E}_{\nu}({q})-\mathbb{E}_{\mu}({p})\geq\varepsilon.$
Theorem 7.
For every pseudometric $d$ we have $\delta^{\mathsf{LP}}(d)=K_{\lambda}(d)=W_{\lambda}(d)$.
Proof.
We only need to show that $K^{\mathsf{rel}}_{\lambda}(d)=K_{\lambda}(d)$, which follows by LABEL:lem:lp-pred-lifting-self-dual and LABEL:lem:self-dual-rel-equals-sym. ∎
5 Convex Powerset Functor
We will next tackle duality for the case of the convex powerset functor, a functor that has been studied in-depth for modelling systems combining probability and non-determinism (e.g., [7]). A non-empty set $D\subseteq\mathcal{D}X$ of probability distributions is convex if for all $\mu_{1},\mu_{2}\in D$ it also holds that $\mu_{1}+_{p}\mu_{2}:=p\cdot\mu_{1}+(1-p)\cdot\mu_{2}\in D$ (where $p\in[0,1]$). For a set $X$, we define
$\mathcal{C}X=\{\emptyset\neq D\subseteq\mathcal{D}X\mid\text{$D$ is convex}\}.$
Of course, we have $\mathcal{C}X\subseteq\mathcal{P}\mathcal{D}X$ for every set $X$. In fact, it is easily verified that every map $\mathcal{P}\mathcal{D}f$ preserves convex sets, so that we obtain a subfunctor $\mathcal{C}$ of the composite functor $\mathcal{P}\mathcal{D}$.
A straightforward – but futile – approach to prove duality for $\mathcal{C}$ would be to observe that it holds for the powerset and distribution functor, and then apply a compositionality result. However, the studied liftings (Kantorovich and Wasserstein lifting) are quite fragile when it comes to compositionality, i.e. it does not hold in general that the composition of liftings of functors $F$, $G$ based on the predicate liftings $\lambda^{F}$, $\lambda^{G}$ is the lifting of the composite $FG$ (based on the obvious combined modality $\lambda^{FG}_{X}=\lambda^{F}_{GX}\circ\lambda^{G}_{X}$) [4]. While it is known that the Wasserstein lifting of the convex powerset functor arises by combining the Wasserstein liftings of the component functors [32], this is incorrect for the Kantorovich lifting [9]. In fact, the given counterexample uses a set that is not convex, thus suggesting that the problem might disappear if we restrict to convex sets.
We use the $\sup$ modality for the powerset functor and the expectation ($\mathbb{E}$) modality for the distribution functor. Our aim is to study the convex powerset functor $\mathcal{C}$ and establish the rather non-trivial result that the combined modality $\lambda_{X}(f)(A)=\sup\{\mathbb{E}_{\mu}({f})\mid\mu\in A\}$ is indeed expressive on its own.
As before, $K_{\lambda}\leq W_{\lambda}$ holds in categorical generality, so the main task is to prove the converse inequality $W_{\lambda}\leq K_{\lambda}$. As discussed earlier, it is known that the Wasserstein representation in terms of $\lambda$ decomposes into the Wasserstein representations in terms of $\sup$ and $\mathbb{E}$, i.e. the Hausdorff and (probabilistic) Kantorovich liftings, respectively:
$W_{\lambda}(d)=W_{\sup}(W_{\mathbb{E}}(d))=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))=:\delta^{\mathsf{HK}}(d)$
To achieve the duality result, it will be convenient to pass from pseudometric spaces to metric spaces. Recall that the metric quotient of a pseudometric space $(X,d)$ is the metric space $(X_{\sim},d_{\sim})$ where $X_{\sim}$ is the set of equivalence classes of the equivalence relation $x\sim y\iff d(x,y)=0$, and $d_{\sim}([x],[y])=d(x,y)$ for any two equivalence classes $x],[y]\in X_{\sim}$.
Lemma 1.
lem:convex-metric-quotient Let $(X,d)$ be a pseudometric space, let $(X_{\sim},d_{\sim})$ be its metric quotient, and let $\pi\colon X\to X_{\sim},x\mapsto[x]$. We then have, for every $A,B\in\mathcal{C}X$,
$K_{\lambda}(d)(A,B)=K_{\lambda}(d_{\sim})(A_{\sim},B_{\sim})\quad\text{and}\quad W_{\lambda}(d)(A,B)=W_{\lambda}(d_{\sim})(A_{\sim},B_{\sim}),$
where $A_{\sim}=\mathcal{C}\pi(A)$ and $B_{\sim}=\mathcal{C}\pi(B)$.
Using the above lemma, we may therefore from now on assume that we are working over a metric space $(X,d)$. We may also assume that $X\neq\emptyset$, as otherwise $\mathcal{C}X=\emptyset$ and both Kantorovich and Wasserstein distance are the empty metric, hence equal.
The main intuition behind the proof is best understood in the case where $X=\{x_{1},\dots,x_{n}\}$ is finite, even though the proof will work for arbitrary $X$. In this case we may view probability distributions and fuzzy predicates on $X$ as vectors in n, and the expectation modality simply computes the dot product between two such vectors: $\mathbb{E}_{\mu}({f})=\mu(x_{1})\cdot f(x_{1})+\dots+\mu(x_{n})\cdot f(x_{n}).$ If $A,B\in\mathcal{C}X$ satisfy $\delta^{\mathsf{HK}}(d)(A,B)>\varepsilon$, then this means, by the definition of the Hausdorff distance, that there must be some $\mu\in A$ such that $\delta^{\mathsf{K}}(\mu,\nu)>\varepsilon$ for all $\nu\in B$ (or we are in the symmetric situation with $A$ and $B$ swapped). This implies that $B$ and the $\varepsilon$-ball around $\mu$ are disjoint convex sets, so we can apply the hyperplane separation theorem to find a hyperplane $H$ such that the two sets lie on opposite sides of that plane. A price function witnessing distance at least $\varepsilon$ under the combined modality $\lambda$ can now be constructed from the normal vector of $H$.
Our proof mostly follows the outline above, but because we may now be working with infinite-dimensional vector spaces, some functional analysis will be required. We leverage this added complexity in Section˜5.2, where we show that the duality result remains true when passing from discrete probability measures to Borel probability measures.
We fix a point $x_{0}\in X$ and consider the vector space $\mathsf{Lip}_{0}(X)$ consisting of the real-valued Lipschitz functions on $X$ vanishing at $x_{0}$:
$\mathsf{Lip}_{0}(X)=\{f\colon X\to\real\mid f(x_{0})=0,\sup_{x\neq y}\textstyle\frac{f(y)-f(x)}{d(x,y)}<\infty\}.$
This is a Banach space with norm given by $\|f\|_{\mathsf{Lip}}=\sup_{x\neq y}\frac{f(y)-f(x)}{d(x,y)}$. We will construct our price function in this space, which is made possible by the fact that the set of probability distributions can be mapped into its dual:
Lemma 2.
lem:embed-into-dual-discrete The set $\mathcal{D}X$ embeds into the continuous dual space $\mathsf{Lip}_{0}(X)^{*}$, that is, every discrete probability measure $\mu$ gives rise to a continuous linear functional $L_{\mu}\colon\mathsf{Lip}_{0}(X)\to\real$, which may explicitly be given by $L_{\mu}(f)=\mathbb{E}_{\mu}({f})$.
In what follows, we often do not distinguish between $\mu$ and $L_{\mu}$ and treat $\mathcal{D}X$ as a subset of $\mathsf{Lip}_{0}(X)^{*}$. We equip $\mathsf{Lip}_{0}(X)^{*}$ with the weak-* topology, which is the weakest topology on $\mathsf{Lip}_{0}(X)^{*}$ making all the maps $\psi\mapsto\psi(f)$ for $f\in\mathsf{Lip}_{0}(X)$ continuous. Equivalently, this is the initial topology wrt. the maps $\psi\mapsto\psi(f)$. Crucially, this topology coincides with the one given by the Kantorovich-Wasserstein distance:
Lemma 3.
lem:topologies-coincide-discrete Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{D}X$ and let $\mu\in\mathcal{D}X$. Then $\mu_{n}\to\mu$ in the topology given by $\delta^{\mathsf{K}}(d)$ iff $L_{\mu_{n}}\to L_{\mu}$ in the weak-* topology.
The space $\mathsf{Lip}_{0}(X)^{*}$ is normed via the operator norm $\|\psi\|_{\mathsf{op}}=\sup_{\|f\|_{\mathsf{Lip}}\leq 1}\psi(f)$. This norm relates to the Wasserstein metric as follows:
Lemma 4.
lem:wasserstein-norm-discrete For any $\mu,\nu\in\mathcal{D}X$ we have $\delta^{\mathsf{K}}(d)(\mu,\nu)=\|\nu-\mu\|_{\mathsf{op}}$.
We are now in a position to state and prove the duality result:
Theorem 5.
thm:convex-powerset-duality For any two convex sets $A,B\in\mathcal{C}X$,
$\delta^{\mathsf{HK}}(d)(A,B)=K_{\lambda}(d)(A,B).$
Proof(sketch).
The inequality $K_{\lambda}\leq\delta^{\mathsf{HK}}$ follows from previous results [4], hence it is sufficient to show $\delta^{\mathsf{HK}}(d)(A,B)\leq K_{\lambda}(d)(A,B)$.
Let $0<\varepsilon<\delta^{\mathsf{HK}}(d)(A,B)$. As outlined above, we may assume wlog. that there exists some $\mu\in A$ such that $\delta^{\mathsf{K}}(d)(\mu,\nu)>\varepsilon$ for every $\nu\in B$. Let $C$ be the closed $\varepsilon$-ball around $\mu$, shifted by $\mu$, and let $D$ be the closure of $B$, also shifted by $\mu$. These two sets are closed and convex, so by the Hahn-Banach separation theorem there exists a continuous linear functional $g\colon\mathsf{Lip}_{0}(X)^{*}\to\real$ and $c\in\real$ such that
$\sup_{\nu\in C}g(\nu)<c\leq\inf_{\nu\in D}g(\nu),$
and because we are in the weak-* topology, the functional $g$ can be represented in the form $\mu\mapsto\mathbb{E}_{\mu}({f})$ for some $f\in\mathsf{Lip}_{0}(X)$. We replace $f$ by $f_{1}=-f/\|f\|_{\mathsf{Lip}}$, which results in a nonexpansive function for which
$|\sup_{\nu\in B}\mathbb{E}_{\nu}({f_{1}})-\sup_{\nu\in A}\mathbb{E}_{\nu}({f_{1}})|\geq\varepsilon.$
The range of $f_{1}$ is not necessarily contained in $[0,1]$, but it must be contained in some subinterval of of length at most $1$ by nonexpansiveness and because $d$ is $1$-bounded. As expectation is linear, we may simply shift $f_{1}$ by a suitable amount. ∎
Remark 2.
A natural question to ask is whether one can, like in the previous section, also obtain a fuzzy-relational version of the duality result. We expect the answer to be ‘yes’, but that it will be necessary to additionally consider the dual predicate lifting $\kappa_{X}(f)(A)=\inf\{\mathbb{E}_{\mu}({f})\mid\mu\in A\}$, resulting in the duality result $K^{\mathsf{rel}}_{\smash{\{\lambda,\kappa\}}}=W_{\lambda}$. This would be reflective of the situation that arises in the case of the Hausdorff extension, where $K^{\mathsf{rel}}_{\{\sup,\inf\}}=W_{\sup}=\delta^{\mathsf{H}}$ holds [32]. We leave this question open for now.
Remark 3(Compositionality).
It has been shown in work on correspondences between Kantorovich and Wasserstein representations of metric liftings [16] (cf. Section 1 under related work) that such correspondences can be combined along sums and product of functors, so that one arrives at general correspondence results for classes of functors obtained by closing given basic building blocks (originally constant functors, identity, powerset, and distributions) under sum and product. One thus obtains correspondences for composite system types such as labelled Markov chains [16, Example 41]. The correspondences produced in this way are not generalized Kantorovich-Rubinstein dualities in the strict sense we use here, as the transition from the Wasserstein presentation to the Kantorovich presentation in general involves the introduction of additional modalities. In particular, this happens for products, where one needs to introduce separate modalities for the factors (indeed, this is what is morally behind the fact that generalized Kantorovich-Rubinstein duality fails for the squaring functor [4]). Nevertheless, our results on generalized Kantorovich-Rubinstein duality for Lévy-Prokhorov distance and convex powerset imply that these functors can now be used as additional basic building blocks in this framework.
5.1 Algorithmic considerations
A nice aspect of the duality result for the convex powerset functor is that it can be used as the basis of an algorithm to compute values of the distance $\delta^{\mathsf{HK}}(d)$. Explicitly, if $(X,d)$ is a finite pseudometric space, and $A_{0},B_{0}\subseteq\mathcal{D}X$ are finite sets of probability measures, the problem is to compute the distance $\delta^{\mathsf{HK}}(d)(A,B)$, where $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$. The distance expands as follows:
$\delta^{\mathsf{HK}}(d)(A,B)=\max(\adjustlimits{\sup}_{\mu\in A}{\inf}_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu),\adjustlimits{\sup}_{\nu\in B}{\inf}_{\mu\in A}\delta^{\mathsf{K}}(d)(\mu,\nu)).$
As the map $\mu\mapsto\inf_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu)$ is convex, the left supremum can instead be taken over $A_{0}$ without changing the value, and similarly we may take the right supremum over $B_{0}$. It is however not in general true that the infima may be taken over $B_{0}$ and $A_{0}$, respectively.
Example 5.
Let $X=\{x,y,z\}$, and assume $d(x,y)=d(x,z)=d(y,z)=1$. Let $A_{0}=\{\mu_{0},\mu_{1}\}$, $B_{0}=\{\mu_{2},\mu_{3}\}$, $\mu_{0}=\frac{1}{3}\cdot x+\frac{1}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{1}=\frac{2}{3}\cdot x+\frac{1}{3}\cdot y$, $\mu_{2}=\frac{2}{3}\cdot y+\frac{1}{3}\cdot z$, $\mu_{3}=\frac{2}{3}\cdot z+\frac{1}{3}\cdot x$. Then the minimal distance from $\mu_{1}$ to $B$ is witnessed by $\mu^{*}=\frac{1}{2}\cdot\mu_{2}+\frac{1}{2}\cdot\mu_{3}=\frac{1}{6}\cdot x+\frac{1}{3}\cdot y+\frac{1}{2}\cdot z$.
Figure 1: Illustration of Example˜5. The thick line segments correspond to the sets $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$, while the shaded region shows the radius-$\frac{1}{2}$ ball around $\mu_{1}$, which is part of a regular hexagon centered at $\mu_{1}$.
To compute the inner infimum, therefore, a more sophisticated approach is required. Çelik et al. [8] consider the problem of solving an optimal transport problem between probability distributions $\mu$ and $\nu$, where $\mu$ is fixed and $\nu$ ranges over some algebraic variety. Their methods can be adapted to our setting, where $\nu$ instead ranges over a convex set $B$:
Consider the typical flow network that one constructs to solve the optimal transport problem, i.e. a complete bipartite graph whose partitions are two copies of the set $X$, which we think of as representing the two probability distributions $\mu$ and $\nu$. A transport plan consists of assigning weights to the edges of this graph in such a way that the sum of weights of incident edges for each vertex matches its probability. The key idea is that there always exists an optimal transport plan that is acyclic in the sense that the edges with positive weight do not contain any cycle. Moreover, if the tree of edges used by the transport plan is known, then the weights of the plan are uniquely determined by the probabilities in $\mu$ and $\nu$ and can be computed by a depth-first (or breadth-first) traversal of the tree. This means that the distance between $\mu$ and $\nu$ can be computed by enumerating all spanning trees of the complete bipartite graph, computing the weights, and taking the least value over all trees where the weights are all non-negative. If $\mu\in A_{0}$ is fixed and $\nu$ ranges over $B$, then the weights in the coupling corresponding to a given spanning tree are linear combinations of the probabilities in $\nu$. The conditions that these weights are non-negative define a linear program whose variables are the coefficients in the convex combination of the elements of $B_{0}$ and whose constraints state that the weights in the tree are all non-negative, and the distance can be found by solving all these linear programs.
The algorithm we just described requires solving exponentially many linear programs in the size of $X$, as there is one such linear program for each spanning tree. If we expand the dual representation of the distance between $A$ and $B$ instead, we obtain
$K_{\lambda}(d)(A,B)=\sup\{|\sup_{\nu\in B}\mathbb{E}_{\nu}({f})-\sup_{\mu\in A}\mathbb{E}_{\mu}({f})|\mid f\colon(X,d)\to(\real,d_{e})\text{ nonexp.}\}.$
This quantity is much easier to compute, as there are no nested suprema/infima. Similar to before, as expectation is linear, the two suprema above may instead be taken over $A_{0}$ and $B_{0}$, respectively. To compute the distance, we can employ the following algorithm. Loop over all pairs $(\mu_{0},\nu_{0})\in A_{0}\times B_{0}$. Given $\mu_{0}$ and $\nu_{0}$, the subproblem is then to find the supremum above for all the nonexpansive $f$ such that $\mathbb{E}_{\mu_{0}}({f})=\sup_{\mu\in A}\mathbb{E}_{\mu}({f})$ and $\mathbb{E}_{\nu_{0}}({f})=\sup_{\nu\in B}\mathbb{E}_{\nu}({f})$. This subproblem can be rephrased as a linear program over the variables $(f_{x})_{x\in X}$ corresponding to the function values of $f$. Nonexpansiveness of $f$, the constraints $\mathbb{E}_{\mu_{0}}({f})\geq\mathbb{E}_{\mu}({f})$ for $\mu\in A_{0}$ and the similar constraints for the $\nu\in B_{0}$ are all easily expressed as linear inequalities between the $f_{x}$. The objective function is $|\mathbb{E}_{\nu_{0}}({f})-\mathbb{E}_{\mu_{0}}({f})|$, which is the maximum of two linear expressions over the $f_{x}$, so we can simply solve the linear program twice, once for each of the two expressions. This algorithm has a runtime complexity that is polynomial in $|X|$, $|A_{0}|$ and $|B_{0}|$, a clear improvement over the exponential complexity for the previous approach.
5.2 Borel Measures
As the categorical Kantorovich and Wasserstein constructions are typically considered in the shape of liftings or lax extensions of set endofunctors, their probabilistic instances are restricted to dealing with discrete probability distributions by necessity. The (probabilistic) Kantorovich-Rubinstein duality, meanwhile, is known to hold for much larger classes of probability distributions, such as Borel measures on metric spaces [30]. In this section we show that this is also true for the convex powerset duality.
For a pseudometric space $(X,d)$ we denote by $\mathsf{Bor}(X,d)$ the set of Borel probability measures, i.e. the probability measures defined on the $\sigma$-algebra generated by the open balls $B^{d}_{\varepsilon}({x})$. Every convex combination of Borel measures is itself a Borel measure. We can therefore also define a functor $\mathcal{C}^{\mathsf{Bor}}\colon\mathbf{PMet}\to\mathbf{PMet}$ where $\mathcal{C}^{\mathsf{Bor}}(X,d)$ is the set of non-empty convex subsets of $\mathsf{Bor}(X,d)$, equipped with Hausdorff-Kantorovich distance $\delta^{\mathsf{HK}}(d)=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))$, where both $\delta^{\mathsf{H}}$ and $\delta^{\mathsf{K}}$ are defined as before. Note that every nonexpansive map on $(X,d)$ is Borel-measurable, so that no issues arise when taking expected values.
$\delta^{\mathsf{K}}(d)(\mu,\nu)=\sup\{|{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}|\mid f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}$
$\delta^{\mathsf{H}}(d)(A,B)=\max(\adjustlimits{\sup}_{x\in A}{\inf}_{y\in B}d(x,y),\adjustlimits{\sup}_{y\in B}{\inf}_{x\in A}d(x,y))$
We emphasize that $\mathcal{C}^{\mathsf{Bor}}$ is not the lifting of a set functor, as the definition of the underlying set depends on the pseudometric on the base space.
Due to a general result by Goncharov et al. [13], we know that $\mathcal{C}^{\mathsf{Bor}}$ admits a Kantorovich representation for some suitable class of predicate liftings. By [13, Theorem 5.3] it suffices to show that it preserves initial morphisms. In the real-valued setting these correspond to isometries, i.e. nonexpansive maps $f\colon(X,d_{1})\to(Y,d_{2})$ such that $d_{2}(f(x),f(x^{\prime}))=d_{1}(x,x^{\prime})$ for all $x,x^{\prime}\in X$.
Lemma 6.
$\mathcal{C}^{\mathsf{Bor}}$ preserves isometries.
The class of predicate liftings one obtains is quite large; we show that one can in fact make do with just a single predicate lifting, which strengthens the corresponding instance of the coalgebraic quantitative Hennessy-Milner theorem [13, Corollary 5.10] by providing a compact explicit syntax for the expressive logic:
Theorem 7.
thm:convex-powerset-duality-borel Let $(X,d)$ be a pseudometric space and $A,B\in\mathcal{C}^{\mathsf{Bor}}(X,d)$. Then
$\delta^{\mathsf{HK}}(d)(A,B)=\sup\{|\sup_{\nu\in B}{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-\sup_{\mu\in A}{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}\,|\mid\\ f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}.$
In the proof, we follow the same steps as before, first passing from pseudometrics to metrics and then leveraging linear algebra to obtain duality.
6 Conclusions and Future Work
We have proved generalized Kantorovich-Rubinstein duality, i.e. coincidence of coupling-based (i.e. Wasserstein) and price-function-based (i.e. Kantorovich or codensity) presentations of functor liftings induced by a given choice of modalities, for two important and non-trivial cases: the Lévy-Prokhorov distance on distributions, and the standard distance on convex sets of distributions that arises from composing the Hausdorff and Kantorovich-Wasserstein metrics. In both cases, we obtain a characterization of the respective distance by means of quantitative modal logics defined by the given modalities; for the case of the Lévy-Prokhorov distances, this logic is (up to restriction of the propositional base) the logic of generally previously studied in context of fuzzy description logics [27], and in the second case the involved modality is just the composite of the usual fuzzy diamond and the expectation modality [32]. In the case of convex powerset, we demonstrate additionally that the Kantorovich presentation plays out algorithmic advantages in the actual computation of distances.
We leave several key open problems, among them on the one hand the extension of the duality result for the Lévy-Prokhorov metric from discrete to Borel probability distributions (while we already prove the duality result for convex powerset at this level of generality), and on the other hand the extension of the duality result for convex powerset to unrestricted fuzzy relations in place of pseudometrics (already established in our result on the Lévy-Prokhorov metric). The latter generalization will amount to a duality result for the known Wasserstein lax extension of convex powerset [32]. Also, we aim to capitalize on the present result in the design of algorithms that actually compute distinguishing formulae as witnesses of lower bounds on behavioural distance, complementing recent results on behavioural distance under the Kantorovich-Wasserstein distance of distributions [24, 28].

### v2

@@
 The classical Kantorovich-Rubinstein duality guarantees coincidence between metrics on the space of probability distributions defined on the one hand via transport plans (couplings) and on the other hand via price functions.
-Both constructions have been lifted to the level of generality of set functors, with the coupling-based construction referred to as the Wasserstein lifting, and the price-function-based construction as the Kantorovich lifting, both based on a choice of quantitative modalities for the given functor.
-It is known that every Wasserstein lifting can be expressed as a Kantorovich lifting; however, the latter in general needs to use additional modalities.
+Both constructions have been lifted to the level of generality of set functors, with the construction based on couplings referred to as the Wasserstein or simply the coupling-based lifting, and the price-function-based construction as the Kantorovich or codensity lifting, both based on a choice of quantitative modalities for the given functor.
+It is known that every coupling-based lifting can be expressed as a price-function-based lifting; however, the latter in general needs to use additional modalities.
 We give an example showing that this cannot be avoided in general.
 We refer to cases in which the same modalities can be used as satisfying the generalized Kantorovich-Rubinstein duality.
-We establish the generalized Kantorovich-Rubinstein duality in this sense for two important cases: The Lévy-Prokhorov distance on distributions, which finds wide-spread applications in machine learning due to its favourable stability properties, and the standard metric on convex sets of distributions that arises by combining the Hausdorff and Kantorovich-Wasserstein distances.
+We establish the generalized Kantorovich-Rubinstein duality in this sense for two important cases: The Lévy-Prokhorov distance on distributions, which finds wide-spread applications in machine learning due to its favourable stability properties, and the standard metric on convex sets of distributions that arises by combining the Hausdorff and Kantorovich-Rubinstein distances.
 1 Introduction
@@
 [36]).
-One well-established metric on the set of distributions over a metric space is variously termed the Kantorovich, Wasserstein, or Hutchinson metric.
-It can be calculated either by minimizing over the expected value of transport plans between, or couplings of, the given distributions, or by maximizing over the difference of expectations taken over all non-expansive price functions.
-The coincidence of these two values is the classical Kantorovich-Rubinstein duality [30, Theorem 5.10].
+One well-established metric on the set of distributions over a metric space is variously termed the Kantorovich-Rubinstein, Wasserstein, or Hutchinson metric.
+It can be calculated either by minimizing over the expected value of transport plans between, or couplings of, the given distributions, or by maximizing over the difference of expectations taken over all nonexpansive price functions.
+The coincidence of these two values is the classical Kantorovich-Rubinstein duality [37, Theorem 5.10].
 Intuitively speaking, a transport plan or coupling is a way to transform one distribution into another by shifting around weight, and its cost, to be minimized, is determined by how much weight is shifted over which distances.
-On the other hand, a price function determines a price for some commodity at given points; non-expansiveness of the price function means that no profit can be made from full-cost transport.
+On the other hand, a price function determines a price for some commodity at given points; nonexpansiveness of the price function means that no profit can be made from full-cost transport.
 The difference between the expected values of a price function under the given distributions is the profit to be made by having the commodity transported, and hence the amount one can offer to a logistics provider when outsourcing the transport.
@@
 In this framework, functors serve as parameters determining a type of systems as their coalgebras; for instance, coalgebras for the distribution functor are Markov chains.
-In the generalized setting, the coupling-based construction is typically referred as the Wasserstein lifting or extension, and the price-function-based one as the Kantorovich lifting or extension.
+In the generalized setting, the coupling-based construction is often referred to as the Wasserstein lifting or extension, and the price-function-based one as the Kantorovich lifting or extension, but other names are found in the literature, such as coupling-based [23] and codensity [25] lifting, respectively.
 Both constructions are parametrized over a choice of quantitative modalities; the classical case involves, on both sides, only one modality, the expectation modality.
-The interest in having a Kantorovich presentation of a given functor lifting or lax extension lies inter alia in the fact that one obtains a quantitative Hennessy-Milner property for the quantitative modal logic generated by the respective modalities [20, 32].
-This property states coincidence of the behavioural distance induced by the given lifting or extension with the logical distance induced by the respective quantitative modal logic; thus, high distance between two states can always be certified by means of a modal formula, a principle that has received increased recent attention (e.g.
-[24, 28]).
+The interest in having a price-function-based presentation of a given functor lifting or lax extension lies inter alia in the fact that one obtains a quantitative Hennessy-Milner property for the quantitative modal logic generated by the respective modalities [27, 39].
+This property states coincidence of the behavioural distance induced by the given lifting or extension with the logical distance induced by the respective quantitative modal logic; thus, high distance between two states can always be certified by means of a modal formula, a prominent principle in the study of behavioural distances (e.g.
+[36, 10, 39, 27, 13] and more recently [31, 35]).
 It has been shown that every metric functor lifting that preserves isometries [14] and every quantitative lax extension [39] is Kantorovich, i.e.
-can be presented via the generalized Kantorovich construction, using however a rather large (in particular typically infinite) set of modalities called the Moss modalities.
-An important open point that remains is thus the question of what we term generalized Kantorovich-Rubinstein duality: In which cases does a Wasserstein distance given by a choice of modalities for a functor coincide with the Kantorovich distance for the same modalities?
-Known positive examples include, as mentioned, the classical Kantorovich-Wasserstein distance on distributions, but also the Hausdorff distance on the powerset.
-We begin our analysis by giving the first example of a natural Wasserstein metric for which generalized Kantorovich-Rubinstein duality in this sense fails, namely the standard $p$-Wasserstein metric for $p>1$ (which minimizes $p$-th roots of the expectation of $p$-th powers of couplings).
+can be presented via the generalized price-function-based construction, using however a rather large (in particular typically infinite) set of modalities called the Moss modalities.
+An important open point that remains is thus the question of what we term generalized Kantorovich-Rubinstein duality: In which cases does a coupling-based distance given by a choice of modalities for a functor coincide with the price-function-based distance for the same modalities?
+Known positive examples include, as mentioned, the classical Kantorovich-Rubinstein distance on distributions, but also the Hausdorff distance on the powerset.
+We begin our analysis by giving an example of a natural coupling-based metric for which generalized Kantorovich-Rubinstein duality in this sense fails, namely the standard $p$-Wasserstein metric for $p>1$ (which minimizes $p$-th roots of the expectation of $p$-th powers of couplings).
 As our main contribution, we then provide two new positive examples, namely the Lévy-Prokhorov distance on probability distributions [30] and the standard distance on the convex powerset, whose elements are convex sets of probability distributions.
-The Lévy-Prokhorov distance has seen a recent rise in popularity due to its favourable robustness properties that make it suitable for tasks in machine learning (such as conformal prediction [3] and corruption resistance [5]); moreover, it has been shown in recent work to induce precisely the behavioural distance defined by $\varepsilon$-bisimilarity [10].
+The Lévy-Prokhorov distance has seen a recent rise in popularity due to its favourable robustness properties that make it suitable for tasks in machine learning (such as conformal prediction [3] and corruption resistance [5]); moreover, it has been shown in recent work [11] to induce precisely the behavioural distance defined by $\varepsilon$-bisimilarity [10].
 In this case, the relevant modality is precisely the generally modality used in work on fuzzy description logics as an alternative formal correspondent of the natural-language term ‘probably’ [34].
 The convex powerset plays a central role in the distribution semantics of Markov decision processes (or probabilistic automata) [7].
-Its standard metric is just the composition of the Hausdorff and Kantorovich-Wasserstein metrics, and as such is again given via the Wasserstein construction for a modality composed from the expectation modality and the standard fuzzy diamond modality [32].
+Its standard metric is just the composition of the Hausdorff and Kantorovich-Rubinstein metrics, and as such is again given via the coupling-based construction for a modality composed from the expectation modality and the standard fuzzy diamond modality [39].
 We show that the generalized Kantorovich-Rubinstein duality holds w.r.t.
 this modality.
-Beyond the mentioned benefits regarding characteristic quantitative modal logics, we demonstrate in this case that the new Kantorovich description also allows for more efficient computation of distances.
+Beyond the mentioned benefits regarding characteristic quantitative modal logics, we demonstrate in this case that the new price-function-based description also allows for more efficient computation of distances.
 Related Work
-Our Kantorovich presentation of convex powerset complements earlier results on a Wasserstein characterization [32] and a quantitative-algebraic presentation [21].
-There has been recent work on what are termed correspondences between the generalized Kantorovich and Wasserstein constructions where a single modality is assumed for the Wasserstein construction, while an associated set of modalities is considered for the Kantorovich construction [16].
+Categorical distance constructions based on couplings were first considered in work on monoidal topology [21, 22].
+The categorical price-function-based construction is predated by several constructions for specific functors in work on stochastic games [17].
+Our price-function-based presentation of convex powerset complements earlier results on a coupling-based characterization [39] and a quantitative-algebraic presentation [28].
+The modality underlying the distance on the convex powerset functor relates to Goubault-Larrecq’s previsions [16]; distances between such previsions are also expressed through a combination of (topological versions of) the Hausdorff and Kantorovich-Rubinstein-distances [17, 18, 19].
+There has been recent work on what are termed correspondences between the generalized coupling-based and price-function-based constructions where a single modality is assumed for the former, while an associated set of modalities is considered for the latter [23].
 Correspondences in this sense thus lie between dualities as considered here, where we insist on the same modalities being used on both sides, and general theorems on Kantorovich presentations of lax extensions [39] and classes of functors [14] that use very large sets of liftings.
-Results are obtained for functors constructed from the main known instances (distributions with the standard Kantorovich-Wasserstein metric, powerset) by applying coproduct and product.
-The Kantorovich construction is sometimes referred to as the codensity construction [18, 19, 16], and as such has been used for logical characterizations of behavioural distances as mentioned above [19] but also for game characterizations [18].
+Results are obtained for functors constructed from the main known instances (distributions with the standard Kantorovich-Rubinstein metric, powerset) by applying coproduct and product.
+The price-function-based construction is sometimes referred to as the codensity construction [25, 26, 23], and as such has been used for logical characterizations of behavioural distances as mentioned above [26] but also for game characterizations [25].
 The problem of generalized Kantorovich-Rubinstein duality has been stated already in work introducing coalgebraic behavioural distances [4], and a simple counterexample has been given; the counterexample we give here is distinguished by involving a quantitative modality that satisfies an analogue of two-valued separation [29].
@@
 Pseudometric spaces and nonexpansive functions between them form a category $\mathbf{PMet}$.
-The full subcategories of $\mathbf{PMet}$ spanned by metric spaces is denoted by $\mathbf{Met}$.
+The full subcategory of $\mathbf{PMet}$ spanned by metric spaces is denoted by $\mathbf{Met}$.
 Coalgebra
 Our main results derive some of their interest from their relevance to behavioural distances in coalgebras.
-Generally, the framework of universal coalgebra [25] is based on abstracting state-based systems as $F$-coalgebras for s functor $F\colon\mathbf{C}\to\mathbf{C}$ on a category $\mathbf{C}$, with $F$ determining the type of the system.
-Specifically, an $F$-coalgebra is a pair $(C,\gamma)$ consisting of a $\mathbf{C}$ object $C$, thought of as an object of states, and a morphism $\gamma\colon C\to FC$ determining transitions from states to structured collections of successor states, with the structure determined by $F$.
+Generally, the framework of universal coalgebra [32] is based on abstracting state-based systems as $F$-coalgebras for a functor $F\colon\mathbf{C}\to\mathbf{C}$ on a category $\mathbf{C}$, with $F$ determining the type of the system.
+Specifically, an $F$-coalgebra is a pair $(C,\gamma)$ consisting of a $\mathbf{C}$-object $C$, thought of as an object of states, and a morphism $\gamma\colon C\to FC$ determining transitions from states to structured collections of successor states, with the structure determined by $F$.
 A homomorphism between $F$-coalgebras $(C,\gamma)$ and $(D,\delta)$ is a $\mathbf{C}$-morphism $h\colon C\to D$ such that $\delta\circ h=Fh\circ\gamma$.
-We list some common functors that occur in the description of $F$ in the definition above.
+We list some common functors that will be useful in the further technical development and their associated coalgebras.
 Example 1.
@@
     The finitely supported probability distribution functor $\mathcal{D}\colon\mathbf{Set}\to\mathbf{Set}$ sends a set $X$ to the set
-    $\mathcal{D}X=\{\mu\colon X\to[0,1]\mid\mu(x)=0\text{ for almost all }x\in X\text{ and }\sum_{x\in X}\mu(x)=1\}$
+    $\mathcal{D}X=\{\mu\colon X\to[0,1]\mid\mu(x)>0\text{ for finitely many }x\in X\text{ and }\sum_{x\in X}\mu(x)=1\}.$
     On a function $f\colon X\to Y$ the functor $\mathcal{D}$ measures probabilities of preimages:
     $\mathcal{D}f(\mu)(y)=\textstyle\sum_{x\in f^{-1}(y)}\mu(x).$
-    The coalgebras of $\mathcal{D}$ are precisely Markov chains.
+    The coalgebras of $\mathcal{D}$ are precisely (discrete-time) Markov chains.
 3 Dual Characterizations of Metrics
@@
   2. 2.
-    The Kantorovich-Wasserstein lifting $\overline{\mathcal{D}}\colon\mathbf{PMet}\to\mathbf{PMet}$ of $\mathcal{D}$ equips the set $\mathcal{D}X$ with the Kantorovich-Wasserstein distance $\delta^{\mathsf{K}}(d_{X})$, which is defined as
-    $\delta^{\mathsf{K}}(d_{X})(\mu,\nu):=\sup\{\mathbb{E}_{\nu}({f})-\mathbb{E}_{\mu}({f})\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexpansive}\},$
+    The Kantorovich-Rubinstein lifting $\overline{\mathcal{D}}\colon\mathbf{PMet}\to\mathbf{PMet}$ of $\mathcal{D}$ equips the set $\mathcal{D}X$ with the Kantorovich-Rubinstein distance $\delta^{\mathsf{KR}}(d_{X})$, which is defined as
+    $\delta^{\mathsf{KR}}(d_{X})(\mu,\nu):=\sup\{\mathbb{E}_{\nu}({f})-\mathbb{E}_{\mu}({f})\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexpansive}\},$
     where $\mathbb{E}_{\mu}({f})=\sum_{x\in X}\mu(x)\cdot f(x)$ denotes the expected value of $f$ under $\mu$.
@@
     Let $\lambda$ be a predicate lifting for $F$.
-The Kantorovich lifting of $F$ to the category of pseudometric spaces sends a metric $d_{X}$ to $K_{\lambda}(d_{X})\colon FX\times FX\to[0,1]$ defined by
+The price-function-based lifting of $F$ to the category of pseudometric spaces sends a metric $d_{X}$ to $K_{\lambda}(d_{X})\colon FX\times FX\to[0,1]$ defined by
     $K_{\lambda}(d_{X})(s,t):=\sup\{|\lambda_{X}(f)(t)-\lambda_{X}(f)(s)|\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexp.}\}$
@@
   3. 3.
-    Let $s,t\in FX$. The set of couplings $\Gamma(s,t)$ is defined as
-    $\Gamma(s,t):=\{c\in F(X\times X)\mid F\pi_{1}(c)=s\text{ and }F\pi_{2}(c)=t\}$
+    Let $s\in FX$ and $t\in FY$. The set of couplings $\Gamma(s,t)$ is defined as
+    $\Gamma(s,t):=\{c\in F(X\times Y)\mid F\pi_{1}(c)=s\text{ and }F\pi_{2}(c)=t\}.$
   4. 4.
     Let $\lambda$ be a well-behaved predicate lifting for $F$ and assume that $F$ preserves weak pullbacks.
-The Wasserstein lifting of $F$ to the category of pseudometric spaces is defined as the lifting that equips $FX$ with $W_{\lambda}(d_{X})$ where
+The coupling-based lifting of $F$ to the category of pseudometric spaces is defined as the lifting that equips $FX$ with $W_{\lambda}(d_{X})$ where
     $W_{\lambda}(d_{X})(s,t)=\inf\{\lambda_{X\times X}(d_{X})(c)\mid c\in\Gamma(s,t)\}.$
-While the Kantorovich lifting assumes no conditions on supplied structures, the Wasserstein lifting is significantly more particular, requiring both pullback preservation of the underlying functor and that the predicate lifting be well-behaved, to ensure that $W_{\lambda}(d_{X})$ is a pseudometric whenever $d_{X}$ is [4, 6].
+While the price-function-based lifting assumes no conditions on supplied structures, the coupling-based lifting is significantly more particular, requiring both pullback preservation of the underlying functor and that the predicate lifting be well-behaved, to ensure that $W_{\lambda}(d_{X})$ is a pseudometric whenever $d_{X}$ is [4, 6].
 Remark 1.
@@
     Let $F=\mathcal{D}$ and let $\lambda$ be the predicate lifting calculating expected values: For $\mu\in\mathcal{D}X$ and $f\in[0,1]^{X}$ we have $\lambda_{X}(f)(\mu)=\mathbb{E}_{\mu}({f})$.
-Then $\delta^{\mathsf{K}}=K_{\lambda}=W_{\lambda}$.
-If, like in the two examples above, the categorical Kantorovich and Wasserstein constructions $K_{\lambda}$ and $W_{\lambda}$ coincide, we say that generalized Kantorovich-Rubinstein duality holds.
+Then $\delta^{\mathsf{KR}}=K_{\lambda}=W_{\lambda}$.
+If, like in the two examples above, the categorical price-function-based and coupling-based constructions $K_{\lambda}$ and $W_{\lambda}$ coincide, we say that generalized Kantorovich-Rubinstein duality holds.
 This name is motivated by the particular case of Example˜3.2, the classical Kantorovich-Rubinstein duality that dates back to the beginnings of transportation theory [24].
@@
 Transportation Theory
-The Wasserstein presentation is closely related to optimal transportation theory: In the case of the Kantorovich-Wasserstein lifting, one may view probability distributions $\mu,\nu$ as producers and consumers of a resource respectively (with the amount of the resource produced/consumed being fixed to $1$).
+The names of the two constructions are motivated by the probabilistic case as well.
+The coupling-based presentation is closely related to optimal transportation theory: In the case of the Kantorovich-Rubinstein lifting, one may view probability distributions $\mu,\nu$ as producers and consumers of a resource respectively (with the amount of the resource produced/consumed being fixed to $1$).
 Now one wants to transport the produced resources to the consumers in the economically most efficient possible way, minimizing the average distance each unit of resource needs to travel.
 Couplings $c$ of $\mu$ and $\nu$ can then be seen as transport plans, with $\mathbb{E}_{c}({d_{X}})$ giving the total cost of carrying out the plan.
-Then the Kantorovich-Wasserstein distance gives us by definition the minimal possible cost.
-In the distribution case (Example˜3.2), the Kantorovich presentation can be explained via the analogy of “outsourcing the transport” by defining a function $f$ that assigns a price to each $x\in X$.
-This function must satisfy the requirement that it is non-expansive, i.e., that the difference of prices assigned to $x,y$ is always at most $d_{X}(x,y)$ (meaning that no extra profit can be made from such a transport).
+Then the Kantorovich-Rubinstein distance gives us by definition the minimal possible cost.
+In the distribution case (Example˜3.2), the price-function-based presentation can be explained via the analogy of “outsourcing the transport” by defining a function $f$ that assigns a price to each $x\in X$.
+This function must satisfy the requirement that it is nonexpansive, i.e., that the difference of prices assigned to $x,y$ is always at most $d_{X}(x,y)$ (meaning that no extra profit can be made from such a transport).
 The overall profit under such a price function is then the income obtained from the consumers ($\mathbb{E}_{\nu}({f})$) minus the cost paid to the producers ($\mathbb{E}_{\mu}({f})$).
-Taking the maximum over all such $f$ gives us the value of the Kantorovich lifting.
+Taking the maximum over all such $f$ gives us the value of the Kantorovich-Rubinstein lifting.
 Expressive Logics
-The Kantorovich lifting is closely related to characteristic multi-valued modal logics: These are logics in which formulae $\varphi$ receive semantics in coalgebras $(X,\gamma)$, inducing an interpretation function $\llbracket\varphi\rrbracket_{\gamma}\colon X\to[0,1]$.
+The price-function-based lifting is closely related to characteristic multi-valued modal logics: These are logics in which formulae $\varphi$ receive semantics in coalgebras $(X,\gamma)$, inducing an interpretation function
+$\llbracket\varphi\rrbracket_{\gamma}\colon X\to[0,1].$
 The logical distance of two states is then the supremum of all distances witnessed by such formulas.
 The semantics of a modal operator $L$ in these types of logics is usually given as a predicate lifting $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$, with $\llbracket L\varphi\rrbracket_{\gamma}$ being inductively defined as $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$.
-Then expressivity of the logic (the fact that the behavioural distance can be witnessed by formulae of the logic arbitrarily closely) can be shown by exploiting the Kantorovich-presentation of the lifting.
+Then expressivity of the logic (the fact that the behavioural distance can be witnessed by formulae of the logic arbitrarily closely) can be shown by exploiting the price-function-based presentation of the lifting.
 In fact, the interpretation $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$ can be viewed as emulating one step of the functional $\Phi_{\gamma}$, under the condition that the interpretations of formulae $\llbracket\varphi\rrbracket_{\gamma}$ are able to approximate any nonexpansive function $f\colon X\to[0,1]$ arbitrarily closely [13].
-Hence, the duality can be used very fruitfully: One can use the Wasserstein view to compute (an under-approximation of) the behavioural distance and switch to Kantorovich to determine the distinguishing formula witnessing this distance.
+Hence, the duality can be used very fruitfully: One can use the coupling-based view to compute (an under-approximation of) the behavioural distance and switch to price functions to determine the distinguishing formula witnessing this distance.
 The $p$-Wasserstein Distance
-There also exists a parametrized version of the Kantorovich-Wasserstein distance, called the $p$-Wasserstein distance for some real parameter $p\geq 1$}.
+There also exists a parametrized version of the Kantorovich-Rubinstein distance, called the $p$-Wasserstein distance for some real parameter $p\geq 1$}.
 In categorical terms, it is given by the predicate lifting $\lambda_{p}(f)(\mu)=(\mathbb{E}_{\mu}({f^{p}}))^{\frac{1}{p}}$ (where $f^{p}$ takes the $p$-th power pointwise), that is:
 $W_{\lambda_{p}}(d)(\mu,\nu)=\inf\{(\mathbb{E}_{\rho}({d^{p}}))^{\frac{1}{p}}\mid\rho\text{ is a coupling of $\mu$ and $\nu$}\}.$
-For $p=1$ this is just the usual Kantorovich-Wasserstein distance and duality holds.
-If $p>1$, however, then duality may fail, and the corresponding Kantorovich construction $K_{\lambda_{p}}$ may be strictly below $W_{\lambda_{p}}$:
+For $p=1$ this is just the usual Kantorovich-Rubinstein distance and duality holds.
+If $p>1$, however, then duality may fail, and the corresponding price-function-based construction $K_{\lambda_{p}}$ may be strictly below $W_{\lambda_{p}}$.
+This gives witness to the general idea from transportation theory that $p=1$ constitutes a special case among the family of Wasserstein distances:
 Example 4.
@@
 We begin by showing the last equality. We note that $d^{2}=d$, so that
-$W_{\lambda_{2}}(d)(\mu,\nu)=(\delta^{\mathsf{K}}(d)(\mu,\nu))^{\frac{1}{2}}=(\tfrac{1}{3})^{\frac{1}{2}}=\tfrac{1}{\sqrt{3}},$
+$W_{\lambda_{2}}(d)(\mu,\nu)=(\delta^{\mathsf{KR}}(d)(\mu,\nu))^{\frac{1}{2}}=(\tfrac{1}{3})^{\frac{1}{2}}=\tfrac{1}{\sqrt{3}},$
 where in the first step we used that $(-)^{\frac{1}{2}}$ is monotone and continuous.
@@
 As the function $g$ is concave, we also have
-$g(0)+g(1)=(\tfrac{1}{3}\cdot g(0)+\tfrac{2}{3}\cdot g(1))+(\tfrac{2}{3}\cdot g(0)+\tfrac{1}{3}\cdot g(1))\\ \leq g(\tfrac{1}{3}\cdot 0+\tfrac{2}{3}\cdot 1)+g(\tfrac{2}{3}\cdot 0+\tfrac{1}{3}\cdot 1)=g(\tfrac{1}{3})+g(\tfrac{2}{3})$ (1)
+$g(0)+g(1)=(\tfrac{1}{3}\cdot g(0)+\tfrac{2}{3}\cdot g(1))+(\tfrac{2}{3}\cdot g(0)+\tfrac{1}{3}\cdot g(1))\\ \leq g(\tfrac{1}{3}\cdot 0+\tfrac{2}{3}\cdot 1)+g(\tfrac{2}{3}\cdot 0+\tfrac{1}{3}\cdot 1)=g(\tfrac{1}{3})+g(\tfrac{2}{3}).$ (1)
+Additionally, nonexpansiveness of $f$ implies that
+$b^{2}-a^{2}=(b-a)(a+b)=(g(1)-g(0))(a+b)\leq(1-0)(a+b)=a+b.$ (2)
 Therefore we have:
@@
 $=\tfrac{1}{3}(b^{2}-a^{2})$
-$=\tfrac{1}{3}(b-a)(a+b)$
-$\leq\tfrac{1}{3}(g(0)+g(1))$ $(\text{$f$ is nonexpansive})$
 […]
-Dividing by $g(\tfrac{2}{3})+g(\tfrac{1}{3})$ on both sides results in the claimed inequality. ∎
+[…]
+Our earlier assumption that $a<b$ implies that $g(\tfrac{2}{3})+g(\tfrac{1}{3})$ is positive, so we can divide by it on both sides, which results in the claimed inequality. ∎
 4 Lévy-Prokhorov Distance
-The Lévy-Prokhorov distance provides an alternative to the Kantorovich-Wasserstein distance when it comes to measuring the distance between probability distributions.
-If $(X,d)$ is pseudometric space, and $\mu,\nu\in\mathcal{D}X$ are (discrete) probability distributions, then we define
+The Lévy-Prokhorov distance provides an alternative to the Kantorovich-Rubinstein distance when it comes to measuring the distance between probability distributions.
+If $(X,d)$ is a pseudometric space, and $\mu,\nu\in\mathcal{D}X$ are (discrete) probability distributions, then we define
 $\delta^{\mathsf{LP}}(d)(\mu,\nu)=\inf\{\varepsilon\mid\forall A\subseteq X.\;\mu(A)\leq\nu(A^{d}_{\varepsilon})+\varepsilon\},$
-where $A^{d}_{\varepsilon}=\{y\in Y\mid\inf_{x\in A}d(x,y)\leq\varepsilon\}$.
-The definition of the Lévy-Prokhorov distance sometimes includes the mirrored condition $\forall B\subseteq Y.\;\nu(B)\leq\mu(B^{d}_{\varepsilon})+\varepsilon$, but this second clause is redundant and does not actually change the induced pseudometric.
+where $A^{d}_{\varepsilon}=\{y\in X\mid\inf_{x\in A}d(x,y)\leq\varepsilon\}$.
+The definition of the Lévy-Prokhorov distance sometimes includes the mirrored condition $\forall B\subseteq X.\;\nu(B)\leq\mu(B^{d}_{\varepsilon})+\varepsilon$, but this second clause is redundant and does not actually change the induced pseudometric.
 The Lévy-Prokhorov distance has recently been investigated by Desharnais and Sokolova [11], who prove that it is a functor lifting, but not a monad lifting, and that it characterizes the notion of $\varepsilon$-bisimulation [10].
-The Lévy-Prokhorov distance admits a Wasserstein representation, which is sometimes also known as the Ky Fan metric.
-It is typically phrased in terms of pairs of (not necessarily independent) random variables that are distributed according to the given distributions, but this is equivalent to considering all their couplings (cf.
-Section˜2).
+The Lévy-Prokhorov distance admits a representation based on couplings (cf.
+Section˜3), or equivalently in terms of pairs of (not necessarily independent) random variables that are distributed according to the given distributions, known as the Ky Fan metric [12].
 The predicate lifting $\lambda$ underlying this representation is given by
@@
 The predicate lifting from (3) has independently been used under the name ‘generally’ in work on fuzzy description logics [34], and it admits a number of equivalent representations.
-Intuitively, all of these representations amount to the statement that the value $\lambda_{X}(f)(\mu)$ is high if the value of $f$ is high with high probability.
+Intuitively, all of these representations amount to the statement that the value $\lambda_{X}(f)(\mu)$ is high if the value of $f$ is high with high probability when sampling according to the distribution $\mu$.
 Lemma 1.
@@
 lem:lp-pred-lifting-self-dual For every $X$, $f$ and $\mu$ we have $\lambda_{X}(f)(\mu)=1-\lambda_{X}(1-f)(\mu)$.
-The Wasserstein representation above is justified by the fact that the predicate lifting is well-behaved:
+The coupling-based representation above is justified by the fact that the predicate lifting is well-behaved:
 Lemma 3.
@@
 Recall that a fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ between sets $X$ and $Y$ is a function $r\colon X\times Y\to[0,1]$.
-The Wasserstein construction applies to fuzzy relations the same way it does to pseudometrics, while the Kantorovich construction is defined in terms of pairs of functions that satisfy a nonexpansiveness condition with respect to the given fuzzy relation.
+The coupling-based construction applies to fuzzy relations the same way it does to pseudometrics, while the price-function-based construction is defined in terms of pairs of functions that satisfy a nonexpansiveness condition with respect to the given fuzzy relation.
 They are therefore both examples of (fuzzy) relational liftings or relators (e.g.
@@
   1. 1.
-    The relational Wasserstein lifting $W_{\lambda}(r)\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
+    The relational coupling-based lifting $W_{\lambda}(r)\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
     $W_{\lambda}(r)(s,t)=\inf\{\lambda_{X\times Y}(r)(c)\mid c\in\Gamma(s,t)\}$
@@
   3. 3.
-    The relational Kantorovich lifting $K^{\mathsf{rel}}_{\lambda}\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
+    The relational price-function-based lifting $K^{\mathsf{rel}}_{\lambda}\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
     $K^{\mathsf{rel}}_{\lambda}(r)(s,t)=\sup\{\lambda_{Y}(g)(t)\ominus\lambda_{X}(f)(s)\mid(f,g)\text{ $r$-nonexpansive}\}$
@@
 lem:self-dual-rel-equals-sym If $\lambda$ is a self-dual predicate lifting, then $K^{\mathsf{rel}}_{\lambda}(d)=K_{\lambda}(d)$ for every pseudometric $d$.
-Out of the two representations of the Lévy-Prokhorov distance discussed earlier, the second, being a Wasserstein lifting, readily generalizes to fuzzy relations.
+Out of the two representations of the Lévy-Prokhorov distance discussed earlier, the second, being based on couplings, readily generalizes to fuzzy relations.
 Therefore, we define the relational Lévy-Prokhorov lifting $\delta^{\mathsf{LP}}$ to be the assignment that maps each fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to $\delta^{\mathsf{LP}}(r)=W_{\lambda}(r)\colon\mathcal{D}X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}\mathcal{D}Y$.
-We discussed in Section˜3 that the inequality $\leq$’ follows from the general theory of Kantorovich and Wasserstein liftings.
-The same is true for the respective lax extensions [31, Lemma 5.22], so that it suffices to prove the converse inequality $\geq$’.
-In the proof of the classical Kantorovich-Rubinstein duality (e.g [30, Theorem 5.10]), this direction amounts to, given an optimal transport plan in the shape of a coupling between the distributions at hand, constructing two price functions that correspond to the optimal cost, in the sense that they form a nonexpansive pair that witnesses the supremum in the definition of the relational Kantorovich lifting $K_{\mathbb{E}}$.
+We discussed in Section˜3 that the inequality $\leq$’ follows from the general theory of coupling-based and price-function-based liftings.
+The same is true for the respective lax extensions [38, Lemma 5.22], so that it suffices to prove the converse inequality $\geq$’.
+In the proof of the classical Kantorovich-Rubinstein duality (e.g [37, Theorem 5.10]), this direction amounts to, given an optimal transport plan in the shape of a coupling between the distributions at hand, constructing two price functions that correspond to the optimal cost, in the sense that they form a nonexpansive pair that witnesses the supremum in the definition of the relational price-function-based lifting $K_{\mathbb{E}}$.
 In our proof of Lévy-Prokhorov duality we use a similar approach, which means that we should first understand how to phrase computation of the Lévy-Prokhorov distance in terms of a transport problem.
 Let $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ and let $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$.
-The Wasserstein representation $\delta^{\mathsf{LP}}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu)$ can be rewritten for this purpose.
+The coupling-based representation $\delta^{\mathsf{LP}}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu)$ can be rewritten for this purpose.
 For $\varepsilon\geq 0$, define $r^{\varepsilon}(x,y)=0$ if $r(x,y)<\varepsilon$ and $r^{\varepsilon}(x,y)=1$ otherwise.
-Then $\rho(\{(x,y)\mid r(x,y)\geq\varepsilon\})=\mathbb{E}_{\rho}({r^{\varepsilon}})$ for every $\rho\in\mathcal{D}(X\times Y)$, and therefore we have
-$W_{\lambda}(r)(\mu,\nu)=\inf\{\inf_{\varepsilon\geq 0}\max(\varepsilon,\mathbb{E}_{\rho}({r^{\varepsilon}}))\mid\rho\in\Gamma(\mu,\nu)\}=\inf_{\varepsilon\geq 0}\max(\varepsilon,W_{\mathbb{E}}(r^{\varepsilon})(\mu,\nu)),$
-where we used the representation of $\lambda$ from LABEL:lem:lp-pred-lifting.1 while replacing the strict inequality by a non-strict one and then swapped the infimum over couplings all the way to the inside.
+Then $\rho(\{(x,y)\mid r(x,y)\geq\varepsilon\})=\mathbb{E}_{\rho}({r^{\varepsilon}})$ for every $\rho\in\mathcal{D}(X\times Y)$.
+We may now use the representation of $\lambda$ from LABEL:lem:lp-pred-lifting.1, replacing the strict inequality with a non-strict one, and then swap the infimum over couplings inside to obtain
+$W_{\lambda}(r)(\mu,\nu)=\inf\{\inf_{\varepsilon\geq 0}\max(\varepsilon,\mathbb{E}_{\rho}({r^{\varepsilon}}))\mid\rho\in\Gamma(\mu,\nu)\}=\inf_{\varepsilon\geq 0}\max(\varepsilon,W_{\mathbb{E}}(r^{\varepsilon})(\mu,\nu)).$
 This means that Lévy-Prokhorov distance is determined by the solutions to the transport problems for the $r^{\varepsilon}$.
@@
 $r^{\varepsilon}$.
-One then replaces the function values 0 and $1$ by $\mathbb{E}_{\mu}({p})$ and $\mathbb{E}_{\mu}({p})+\varepsilon$ to arrive at a $d$-nonexpansive pair $(f,g)$.
+One then replaces the function values 0 and $1$ by $\mathbb{E}_{\mu}({p})$ and $\mathbb{E}_{\mu}({p})+\varepsilon$ to arrive at an $r$-nonexpansive pair $(f,g)$.
 Using this pair we show
-$K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)\geq\lambda_{Y}(g)(\nu)-\lambda_{X}(f)(\mu)\geq\mathbb{E}_{\nu}({q})-\mathbb{E}_{\mu}({p})\geq\varepsilon.$
+$K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)\geq\lambda_{Y}(g)(\nu)-\lambda_{X}(f)(\mu)\geq(\mathbb{E}_{\mu}({p})+\varepsilon)-\mathbb{E}_{\mu}({p})=\varepsilon.$
 Theorem 7.
@@
 A straightforward – but futile – approach to prove duality for $\mathcal{C}$ would be to observe that it holds for the powerset and distribution functor, and then apply a compositionality result.
-However, the studied liftings (Kantorovich and Wasserstein lifting) are quite fragile when it comes to compositionality, i.e.
+However, the studied liftings (price-function-based and coupling-based) are quite fragile when it comes to compositionality, i.e.
 it does not hold in general that the composition of liftings of functors $F$, $G$ based on the predicate liftings $\lambda^{F}$, $\lambda^{G}$ is the lifting of the composite $FG$ (based on the obvious combined modality $\lambda^{FG}_{X}=\lambda^{F}_{GX}\circ\lambda^{G}_{X}$) [4].
-While it is known that the Wasserstein lifting of the convex powerset functor arises by combining the Wasserstein liftings of the component functors [32], this is incorrect for the Kantorovich lifting [9].
+While it is known that the coupling-based lifting of the convex powerset functor arises by combining the coupling-based liftings of the component functors [39], this is incorrect for the price-function-based lifting [9].
 In fact, the given counterexample uses a set that is not convex, thus suggesting that the problem might disappear if we restrict to convex sets.
@@
 As before, $K_{\lambda}\leq W_{\lambda}$ holds in categorical generality, so the main task is to prove the converse inequality $W_{\lambda}\leq K_{\lambda}$.
-As discussed earlier, it is known that the Wasserstein representation in terms of $\lambda$ decomposes into the Wasserstein representations in terms of $\sup$ and $\mathbb{E}$, i.e.
-the Hausdorff and (probabilistic) Kantorovich liftings, respectively:
-$W_{\lambda}(d)=W_{\sup}(W_{\mathbb{E}}(d))=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))=:\delta^{\mathsf{HK}}(d)$
+As discussed earlier, it is known that the coupling-based representation in terms of $\lambda$ decomposes into the coupling-based representations in terms of $\sup$ and $\mathbb{E}$, i.e.
+the Hausdorff and Kantorovich-Rubinstein liftings, respectively:
+$W_{\lambda}(d)=W_{\sup}(W_{\mathbb{E}}(d))=\delta^{\mathsf{H}}(\delta^{\mathsf{KR}}(d))=:\delta^{\mathsf{HK}}(d)$
 To achieve the duality result, it will be convenient to pass from pseudometric spaces to metric spaces.
@@
 Using the above lemma, we may therefore from now on assume that we are working over a metric space $(X,d)$.
-We may also assume that $X\neq\emptyset$, as otherwise $\mathcal{C}X=\emptyset$ and both Kantorovich and Wasserstein distance are the empty metric, hence equal.
+We may also assume that $X\neq\emptyset$, as otherwise $\mathcal{C}X=\emptyset$ and both coupling-based and price-function-based distance are the empty metric, hence equal.
 The main intuition behind the proof is best understood in the case where $X=\{x_{1},\dots,x_{n}\}$ is finite, even though the proof will work for arbitrary $X$.
-In this case we may view probability distributions and fuzzy predicates on $X$ as vectors in n, and the expectation modality simply computes the dot product between two such vectors: $\mathbb{E}_{\mu}({f})=\mu(x_{1})\cdot f(x_{1})+\dots+\mu(x_{n})\cdot f(x_{n}).$ If $A,B\in\mathcal{C}X$ satisfy $\delta^{\mathsf{HK}}(d)(A,B)>\varepsilon$, then this means, by the definition of the Hausdorff distance, that there must be some $\mu\in A$ such that $\delta^{\mathsf{K}}(\mu,\nu)>\varepsilon$ for all $\nu\in B$ (or we are in the symmetric situation with $A$ and $B$ swapped).
-This implies that $B$ and the $\varepsilon$-ball around $\mu$ are disjoint convex sets, so we can apply the hyperplane separation theorem to find a hyperplane $H$ such that the two sets lie on opposite sides of that plane.
+In this case we may view probability distributions and fuzzy predicates on $X$ as vectors in n, and the expectation modality simply computes the dot product between two such vectors: $\mathbb{E}_{\mu}({f})=\mu(x_{1})\cdot f(x_{1})+\dots+\mu(x_{n})\cdot f(x_{n}).$ If $A,B\in\mathcal{C}X$ satisfy $\delta^{\mathsf{HK}}(d)(A,B)>\varepsilon$, then this means, by the definition of the Hausdorff distance, that there must be some $\mu\in A$ such that $\delta^{\mathsf{KR}}(\mu,\nu)>\varepsilon$ for all $\nu\in B$ (or we are in the symmetric situation with $A$ and $B$ swapped).
+This implies that $B$ and the $\varepsilon$-ball around $\mu$ are disjoint convex sets, so we can apply the hyperplane separation theorem to find a hyperplane $H$ such that the two sets lie on opposite sides of that plane (more precisely, $B$ may intersect the hyperplane, but the $\varepsilon$-ball may not).
 A price function witnessing distance at least $\varepsilon$ under the combined modality $\lambda$ can now be constructed from the normal vector of $H$.
@@
 the maps $\psi\mapsto\psi(f)$.
-Crucially, this topology coincides with the one given by the Kantorovich-Wasserstein distance:
+Crucially, this topology coincides with the one given by the Kantorovich-Rubinstein distance:
 Lemma 3.
 lem:topologies-coincide-discrete Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{D}X$ and let $\mu\in\mathcal{D}X$.
-Then $\mu_{n}\to\mu$ in the topology given by $\delta^{\mathsf{K}}(d)$ iff $L_{\mu_{n}}\to L_{\mu}$ in the weak-* topology.
-The space $\mathsf{Lip}_{0}(X)^{*}$ is normed via the operator norm $\|\psi\|_{\mathsf{op}}=\sup_{\|f\|_{\mathsf{Lip}}\leq 1}\psi(f)$. This norm relates to the Wasserstein metric as follows:
+Then $\mu_{n}\to\mu$ in the topology given by $\delta^{\mathsf{KR}}(d)$ iff $L_{\mu_{n}}\to L_{\mu}$ in the weak-* topology.
+The space $\mathsf{Lip}_{0}(X)^{*}$ is normed via the operator norm $\|\psi\|_{\mathsf{op}}=\sup_{\|f\|_{\mathsf{Lip}}\leq 1}\psi(f)$.
+This norm relates to the Kantorovich-Rubinstein metric as follows:
 Lemma 4.
-lem:wasserstein-norm-discrete For any $\mu,\nu\in\mathcal{D}X$ we have $\delta^{\mathsf{K}}(d)(\mu,\nu)=\|\nu-\mu\|_{\mathsf{op}}$.
+lem:wasserstein-norm-discrete For any $\mu,\nu\in\mathcal{D}X$ we have $\delta^{\mathsf{KR}}(d)(\mu,\nu)=\|\nu-\mu\|_{\mathsf{op}}$.
 We are now in a position to state and prove the duality result:
@@
 As outlined above, we may assume wlog.
-that there exists some $\mu\in A$ such that $\delta^{\mathsf{K}}(d)(\mu,\nu)>\varepsilon$ for every $\nu\in B$.
+that there exists some $\mu\in A$ such that $\delta^{\mathsf{KR}}(d)(\mu,\nu)>\varepsilon$ for every $\nu\in B$.
 Let $C$ be the closed $\varepsilon$-ball around $\mu$, shifted by $\mu$, and let $D$ be the closure of $B$, also shifted by $\mu$.
@@
 The range of $f_{1}$ is not necessarily contained in $[0,1]$, but it must be contained in some subinterval of of length at most $1$ by nonexpansiveness and because $d$ is $1$-bounded.
-As expectation is linear, we may simply shift $f_{1}$ by a suitable amount.
+As expectation is linear, we may simply shift $f_{1}$ by a suitable amount to extract the desired price function.
 ∎
@@
 Remark 3(Compositionality).
-It has been shown in work on correspondences between Kantorovich and Wasserstein representations of metric liftings [16] (cf.
+It has been shown in work on correspondences between price-function-based and coupling-based representations of metric liftings [23] (cf.
 Section 1 under related work) that such correspondences can be combined along sums and product of functors, so that one arrives at general correspondence results for classes of functors obtained by closing given basic building blocks (originally constant functors, identity, powerset, and distributions) under sum and product.
-One thus obtains correspondences for composite system types such as labelled Markov chains [16, Example 41].
-The correspondences produced in this way are not generalized Kantorovich-Rubinstein dualities in the strict sense we use here, as the transition from the Wasserstein presentation to the Kantorovich presentation in general involves the introduction of additional modalities.
+One thus obtains correspondences for composite system types such as labelled Markov chains [23, Example 41].
+The correspondences produced in this way are not generalized Kantorovich-Rubinstein dualities in the strict sense we use here, as the transition from the coupling-based presentation to the price-function-based presentation in general involves the introduction of additional modalities.
 In particular, this happens for products, where one needs to introduce separate modalities for the factors (indeed, this is what is morally behind the fact that generalized Kantorovich-Rubinstein duality fails for the squaring functor [4]).
@@
 The distance expands as follows:
-$\delta^{\mathsf{HK}}(d)(A,B)=\max(\adjustlimits{\sup}_{\mu\in A}{\inf}_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu),\adjustlimits{\sup}_{\nu\in B}{\inf}_{\mu\in A}\delta^{\mathsf{K}}(d)(\mu,\nu)).$
-As the map $\mu\mapsto\inf_{\nu\in B}\delta^{\mathsf{K}}(d)(\mu,\nu)$ is convex, the left supremum can instead be taken over $A_{0}$ without changing the value, and similarly we may take the right supremum over $B_{0}$.
+$\delta^{\mathsf{HK}}(d)(A,B)=\max(\adjustlimits{\sup}_{\mu\in A}{\inf}_{\nu\in B}\delta^{\mathsf{KR}}(d)(\mu,\nu),\adjustlimits{\sup}_{\nu\in B}{\inf}_{\mu\in A}\delta^{\mathsf{KR}}(d)(\mu,\nu)).$
+As the map $\mu\mapsto\inf_{\nu\in B}\delta^{\mathsf{KR}}(d)(\mu,\nu)$ is convex, the left supremum can instead be taken over $A_{0}$ without changing the value, and similarly we may take the right supremum over $B_{0}$.
 It is however not in general true that the infima may be taken over $B_{0}$ and $A_{0}$, respectively.
@@
 Then the minimal distance from $\mu_{1}$ to $B$ is witnessed by $\mu^{*}=\frac{1}{2}\cdot\mu_{2}+\frac{1}{2}\cdot\mu_{3}=\frac{1}{6}\cdot x+\frac{1}{3}\cdot y+\frac{1}{2}\cdot z$.
+See Figure˜1 for an illustration.
 Figure 1: Illustration of Example˜5.
-The thick line segments correspond to the sets $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$, while the shaded region shows the radius-$\frac{1}{2}$ ball around $\mu_{1}$, which is part of a regular hexagon centered at $\mu_{1}$.
+The thick line segments correspond to the sets $A=\mathsf{conv}(A_{0})$ and $B=\mathsf{conv}(B_{0})$, while the shaded region shows the radius-$\frac{1}{2}$-ball around $\mu_{1}$, which is part of a regular hexagon centered at $\mu_{1}$.
 To compute the inner infimum, therefore, a more sophisticated approach is required.
@@
 5.2 Borel Measures
-As the categorical Kantorovich and Wasserstein constructions are typically considered in the shape of liftings or lax extensions of set endofunctors, their probabilistic instances are restricted to dealing with discrete probability distributions by necessity.
-The (probabilistic) Kantorovich-Rubinstein duality, meanwhile, is known to hold for much larger classes of probability distributions, such as Borel measures on metric spaces [30].
+As the categorical coupling-based and price-function-based constructions are typically considered in the shape of liftings or lax extensions of set endofunctors, their probabilistic instances are restricted to dealing with discrete probability distributions by necessity.
+The (probabilistic) Kantorovich-Rubinstein duality, meanwhile, is known to hold for much larger classes of probability distributions, such as Radon measures on metric spaces [37].
 In this section we show that this is also true for the convex powerset duality.
@@
 Every convex combination of Borel measures is itself a Borel measure.
-We can therefore also define a functor $\mathcal{C}^{\mathsf{Bor}}\colon\mathbf{PMet}\to\mathbf{PMet}$ where $\mathcal{C}^{\mathsf{Bor}}(X,d)$ is the set of non-empty convex subsets of $\mathsf{Bor}(X,d)$, equipped with Hausdorff-Kantorovich distance $\delta^{\mathsf{HK}}(d)=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))$, where both $\delta^{\mathsf{H}}$ and $\delta^{\mathsf{K}}$ are defined as before.
+We can therefore also define a functor $\mathcal{C}^{\mathsf{Bor}}\colon\mathbf{PMet}\to\mathbf{PMet}$ where $\mathcal{C}^{\mathsf{Bor}}(X,d)$ is the set of non-empty convex subsets of $\mathsf{Bor}(X,d)$, equipped with Hausdorff-Kantorovich distance $\delta^{\mathsf{HK}}(d)=\delta^{\mathsf{H}}(\delta^{\mathsf{KR}}(d))$, where both $\delta^{\mathsf{H}}$ and $\delta^{\mathsf{KR}}$ are defined as before.
 Note that every nonexpansive map on $(X,d)$ is Borel-measurable, so that no issues arise when taking expected values.
-$\delta^{\mathsf{K}}(d)(\mu,\nu)=\sup\{|{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}|\mid f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}$
+$\delta^{\mathsf{KR}}(d)(\mu,\nu)=\sup\{|{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}|\mid f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}$
 $\delta^{\mathsf{H}}(d)(A,B)=\max(\adjustlimits{\sup}_{x\in A}{\inf}_{y\in B}d(x,y),\adjustlimits{\sup}_{y\in B}{\inf}_{x\in A}d(x,y))$
@@
 [14], we know that $\mathcal{C}^{\mathsf{Bor}}$ admits a Kantorovich representation for some suitable class of predicate liftings.
-By [13, Theorem 5.3] it suffices to show that it preserves initial morphisms.
+By [14, Theorem 5.3] it suffices to show that it preserves initial morphisms.
 In the real-valued setting these correspond to isometries, i.e.
@@
 $\mathcal{C}^{\mathsf{Bor}}$ preserves isometries.
-The class of predicate liftings one obtains is quite large; we show that one can in fact make do with just a single predicate lifting, which strengthens the corresponding instance of the coalgebraic quantitative Hennessy-Milner theorem [13, Corollary 5.10] by providing a compact explicit syntax for the expressive logic:
+The class of predicate liftings one obtains is quite large; we show that one can in fact make do with just a single predicate lifting, which strengthens the corresponding instance of the coalgebraic quantitative Hennessy-Milner theorem [14, Corollary 5.10] by providing a compact explicit syntax for the expressive logic:
 Theorem 7.
@@
 $\delta^{\mathsf{HK}}(d)(A,B)=\sup\{|\sup_{\nu\in B}{\textstyle\int_{X}{f}\,\mathrm{d}{\nu}}-\sup_{\mu\in A}{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}\,|\mid\\ f\colon(X,d)\to([0,1],d_{e})\text{ nonexp.}\}.$
-In the proof, we follow the same steps as before, first passing from pseudometrics to metrics and then leveraging linear algebra to obtain duality.
+In the proof, we follow the same steps as before, first passing from pseudometrics to metrics and then leveraging linear algebra to obtain coincidence of the two distances.
+Remark 4.
+LABEL:thm:convex-powerset-duality-borel establishes that the composition of the Hausdorff and Kantorovich-Rubinstein distances can be expressed in price-function form with respect to the single modality $A\mapsto\sup_{\mu\in A}{\textstyle\int_{X}{f}\,\mathrm{d}{\mu}}$.
+To extend this into a full-blown duality result for continuous measures, stating that $\delta^{\mathsf{HK}}$ also coincides with the corresponding coupling-based construction, two additional steps are needed: First, we need to have duality for the two individual distances.
+In the case of the Hausdorff distance, no additional requirements are necessary, but in the case of Kantorovich-Rubinstein distance one needs to restrict from Borel measures over general pseudometric spaces to one of the settings in which duality is known to hold, such as Polish spaces or Radon measures [37].
+Second, one needs to establish a compositionality result for the respective coupling-based constructions, generalizing the one from the discrete case [39, Example 6.8.3].
+We leave the details of such a duality result for future work.
 6 Conclusions and Future Work
 We have proved generalized Kantorovich-Rubinstein duality, i.e.
-coincidence of coupling-based (i.e.
-Wasserstein) and price-function-based (i.e.
-Kantorovich or codensity) presentations of functor liftings induced by a given choice of modalities, for two important and non-trivial cases: the Lévy-Prokhorov distance on distributions, and the standard distance on convex sets of distributions that arises from composing the Hausdorff and Kantorovich-Wasserstein metrics.
+coincidence of coupling-based (or Wasserstein) and price-function-based (or Kantorovich or codensity) presentations of functor liftings induced by a given choice of modalities, for two important and non-trivial cases: the Lévy-Prokhorov distance on distributions, and the standard distance on convex sets of distributions that arises from composing the Hausdorff and Kantorovich-Rubinstein metrics.
 In both cases, we obtain a characterization of the respective distance by means of quantitative modal logics defined by the given modalities; for the case of the Lévy-Prokhorov distances, this logic is (up to restriction of the propositional base) the logic of generally previously studied in context of fuzzy description logics [34], and in the second case the involved modality is just the composite of the usual fuzzy diamond and the expectation modality [39].
-In the case of convex powerset, we demonstrate additionally that the Kantorovich presentation plays out algorithmic advantages in the actual computation of distances.
-We leave several key open problems, among them on the one hand the extension of the duality result for the Lévy-Prokhorov metric from discrete to Borel probability distributions (while we already prove the duality result for convex powerset at this level of generality), and on the other hand the extension of the duality result for convex powerset to unrestricted fuzzy relations in place of pseudometrics (already established in our result on the Lévy-Prokhorov metric).
-The latter generalization will amount to a duality result for the known Wasserstein lax extension of convex powerset [32].
-Also, we aim to capitalize on the present result in the design of algorithms that actually compute distinguishing formulae as witnesses of lower bounds on behavioural distance, complementing recent results on behavioural distance under the Kantorovich-Wasserstein distance of distributions [24, 28].
+In the case of convex powerset, we demonstrate additionally that the price-function-based presentation plays out algorithmic advantages in the actual computation of distances.
+We leave several key open problems, among them on the one hand the extension of the duality result for the Lévy-Prokhorov metric from discrete to Borel probability distributions, and on the other hand the extension of the duality result for convex powerset to unrestricted fuzzy relations in place of pseudometrics (already established in our result on the Lévy-Prokhorov metric).
+The latter generalization will amount to a duality result for the known coupling-based lax extension of convex powerset [39].
+Also, we aim to capitalize on the present result in the design of algorithms that actually compute distinguishing formulae as witnesses of lower bounds on behavioural distance, complementing recent results on behavioural distance under the Kantorovich-Rubinstein distance of distributions [31, 35].